### v1

# N-Agent Ad Hoc Teamwork

Abstract
Current approaches to learning cooperative behaviors in multi-agent settings assume relatively restrictive settings. In standard fully cooperative multi-agent reinforcement learning, the learning algorithm controls all agents in the scenario, while in ad hoc teamwork, the learning algorithm usually assumes control over only a single agent in the scenario. However, many cooperative settings in the real world are much less restrictive. For example, in an autonomous driving scenario, a company might train its cars with the same learning algorithm, yet once on the road, these cars must cooperate with cars from another company. Towards generalizing the class of scenarios that cooperative learning methods can address, we introduce $N$-agent ad hoc teamwork, in which a set of autonomous agents must interact and cooperate with dynamically varying numbers and types of teammates at evaluation time. This paper formalizes the problem, and proposes the Policy Optimization with Agent Modelling (POAM) algorithm. POAM is a policy gradient, multi-agent reinforcement learning approach to the NAHT problem, that enables adaptation to diverse teammate behaviors by learning representations of teammate behaviors. Empirical evaluation on StarCraft II tasks shows that POAM improves cooperative task returns compared to baseline approaches, and enables out-of-distribution generalization to unseen teammates.
1 Introduction
Cooperative multi-agent reinforcement learning (MARL) is a paradigm for learning agent teams that solve a common task via interaction with each other and the environment (Littman, 1994). A related paradigm for learning cooperative behavior is ad hoc teamwork (AHT). In contrast to MARL, the objective of AHT is to create a single agent policy that can collaborate with unknown teammates to solve a common task (Mirsky et al., 2022; Stone et al., 2010).
While impressive, these advances have largely been confined to scenarios in which either complete control over all agents is assumed, or only a single agent is adapted for cooperation (Rashid et al., 2018; Son et al., 2019; Lin et al., 2023; Baker et al., 2019). However, real-world collaborative scenarios—e.g. search-and-rescue, or robot fleets for warehouses—might demand agent subteams that are able to collaborate with unfamiliar teammates that follow different coordination conventions.
Towards producing agent teams that are more flexible and applicable to realistic cooperative scenarios, this paper proposes the novel problem setting of $N$-agent ad hoc teamwork (NAHT), in which a set of autonomous agents must interact with an uncontrolled set of teammates to perform a cooperative task. When there is only a single ad hoc agent, NAHT is equivalent to AHT. On the other hand, when all ad hoc agents were produced by some MARL algorithm and there are no uncontrolled teammates, NAHT is equivalent to MARL. Thus, the proposed problem setting generalizes both MARL and AHT.
Drawing from ideas in both MARL and AHT, we introduce Policy Optimization with Agent Modelling (POAM). POAM is a policy-gradient based approach for learning cooperative multi-agent team behaviors, in the presence of varying numbers and types of teammate behaviors. It consists of (1) an agent modeling network that generates a vector characterizing teammate behaviors, and (2) an independent actor-critic architecture, which conditions on the learned teammate vectors to enable adaptation to a variety of potential teammate behaviors. Empirical evaluation on StarCraft II tasks shows that POAM learns to coordinate with a changing number of teammates of various types, with higher competency than MARL, AHT, and NAHT baseline approaches. An evaluation with out-of-distribution teammates also reveals that POAM’s agent modeling module enables improved generalization to out-of-distribution teammates, compared to baseline without agent modeling.
To summarize, this paper makes the following contributions:
  • It proposes, motivates, and formalizes the novel problem setting of N-agent ad hoc teamwork.
  • It introduces Policy Optimization with Agent Modelling (POAM), an algorithm that leverages ideas from AHT and MARL to enable learning cooperative team behaviors whilst dealing with diverse behaviors from uncontrolled teammates.
  • It empirically demonstrates that POAM enables improved cooperative task returns over baselines, and generalization to out-of-distribution teammates.
Figure 1: Left: MARL algorithms assume full control over all $M$ agents in a cooperative scenario. Center: AHT algorithms often assume that only a single agent is controlled by the learning algorithm, while the other $M-1$ agents are uncontrolled and can have a diverse, unknown set of behaviors. Right: NAHT, the paradigm proposed by this paper, assumes that $N$ agents (where $N$ can vary) are controlled by the learning algorithm, while the remaining $M-N$ agents are uncontrolled.
2 Background
[…]
3 NAHT Problem Formulation
Drawing from the goals of MARL and AHT (Stone et al., 2010), the goal of $N$-agent ad-hoc teamwork is to create a set of autonomous agents that are able to efficiently collaborate with both known and unknown teammates to maximize return on a task. The goal is formalized below.
Let $C$ denote a set of ad-hoc agents. If the policies of the agents in $C$ are generated by an algorithm, we say that the algorithm controls agents in $C$. Since our intention is to develop algorithms for generating the policies of agents in $C$, we refer to agents in $C$ as controlled. Let $U$ denote a set of uncontrolled agents, which we define as all agents in the environment not included in $C$ [^3] The original AHT problem statement used the terms “known” and “unknown” agents. However, it is common for modern AHT learning methods to assume some knowledge about the “unknown” teammates during training (Mirsky et al., 2022). Thus, we instead employ the terms controlled and uncontrolled. We adopt an analogous assumption to Stone et al. (2010)—namely, that agents in $U$ satisfy a minimal level of competency in a team setting.
We model an open system of interaction, in which a random selection of $M$ agents from sets $C$ and $U$ must coordinate to perform task $r$. For illustration, consider a warehouse staffed by robots developed by companies $A$ and $B$, where there is a box-lifting task that requires three robots to accomplish. If Company A’s robots are ad-hoc agents (corresponding to $C$), then some robots from Company A could collaborate with robots from Company B (corresponding to $U$) to accomplish the task, rather than requiring that all three robots come exclusively from $A$ or $B$. Motivated thus, we introduce a team sampling procedure $X$. At the beginning of each episode, $X$ samples a team of $M$ agents by first sampling $N<M$, then sampling $N$ agents from the set $C$ and $M-N$ agents from $U$. We define $X$ more formally below.
Let $g(k,\mu,s)$ represent a function for randomly selecting $k$ elements from a generic set $\mu$, conditioned on a state $s\in\mathcal{S}$. For instance, $g$ might be the distribution, Multinomial$(k,|\mu|,\bm{\phi}(s))$, for selecting $k$ elements with replacement from the set $\mu$, parameterized by the state-dependent probability logits $\bm{\phi}(s)\in[0,1]^{|\mu|}$. Allowing $g$ to depend on an initial state enriches the representable open interactions. Returning to our robot warehouse example, a state-dependent $g$ might enable us to model that robots suitable for heavy loads are more likely to be present near the loading dock area. For $s\in\mathcal{S}$, X is a sampling procedure parameterized by the tuple $(s,M,U,C,\bm{\phi}_{M},g_{U},g_{C})$, where $\bm{\phi}_{M}\in[0,1]^{M}$ represents the logits of a categorical distribution:
  1. 1.
    Sample an integer $N$ from Cat$(\bm{\phi}_{M})$.
  2. 2.
    Sample $N$ controlled agents via $g_{U}$ from $U$.
  3. 3.
    Sample $M-N$ uncontrolled agents via $g_{C}$ from $C$.
Since teams consisting entirely of unknown agents are not of interest, we restrict consideration to teams containing at least one controlled agent, i.e $N\geq 1$. The agent sampling functions $g_{U},g_{C}$ could employ sampling with replacement to model scenarios where two robots may use the same policy, or sampling without replacement for scenarios where that is not possible (e.g. if all robots are heterogeneous).
Without loss of generality, let $C(\theta)$ denote a set of agents parameterized by $\theta$. Let the notation $\bm{\pi}^{(M)}\sim X(U,C)$ indicate sampling a team of size $M$ from $U$ and $C$. The objective of the NAHT problem is to find parameters $\theta$, such that $C(\theta)$ maximizes the expected return in the presence of teammates from $U$:
$\max_{\theta}\left(\mathbb{E}_{\bm{\pi}^{(M)}\sim X(U,C(\theta))}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}\right]\right).$ (1)
The central challenges of the NAHT problem include (1) coordinating with potentially unknown teammates (generalization), and (2) coping with an unknown number of uncontrolled teammates (openness).
4 A Motivating Example
Having introduced the NAHT problem, a natural question to consider is whether AHT solutions may optimally address NAHT problems as well. If so, then there would be little need to consider the NAHT problem setting. For instance, as an AHT solution consists of a single teammate policy, trained in the scenario illustrated in Figure 1, a simple yet reasonable approach consists of directly copying the AHT policy as many times as required, in an NAHT problem.
To illustrate the limitations of such an approach, we consider a concrete example in which a single ad hoc agent would follow a rational policy, but two ad hoc agents following that same policy would lead to a suboptimal outcome. Define the following simple game for $M$ agents: at each turn, each agent $a_{i}$ picks one bit $b_{i}\in\{0,1\}$; at the end of each turn, all the bits are summed $s=\sum_{i}b_{i}$. The team wins if and only if the sum of the chosen bits is exactly $1$.
Consider a scenario where there are $M-1$ (uncontrolled) agents who follow a deterministic policy. For simplicity, suppose they always pick 00. We now introduce a new ad hoc agent $a_{h}$. Assume that the ad hoc agent observes the $M-1$ always-zero agents play for one step, and the ad hoc agent $a_{h}$ is introduced into the game in the second step. Suppose the ad hoc agent selects actions according to the following strategy:
  1. 1.
    Model teammate actions by observing the last step.
  2. 2.
    Select the best response action.
  3. 3.
    If no information is known about a teammate, assume they act according to the average of expected actions of all the other observed agents.
If the ad hoc agent $a_{h}$ acts based on the teammate modeling scheme described above, it will observe that all the other agents picked 00 at the first step, and therefore pick $1$.
Now, consider a scenario in which there are $M-2$ always-zero agents playing without the ad hoc agent for one step, and the ad hoc agents $a_{h}^{1}$ and $a_{h}^{2}$ are introduced into the game in the second step. Following the same modeling policy, the two ad hoc agents $a_{h}^{1}$ and $a_{h}^{2}$ will both assume that their teammates follow an always-zero policy, and therefore learn to pick the 1 action—a suboptimal decision, since $s=2$. Following the second step, both $a_{h}^{1}$ and $a_{h}^{2}$ will assume that their counterparts pick the $1$ action. As a result, both agents will switch to the action they would consider the best response, and would both pick the 00 action, meaning once again the joint action is not optimal, as $s=0$. We observe that extending the agent modeling window beyond one step does not resolve this inherent pattern (see Appendix A.4).
While the policy above leads to optimal behavior for a single ad hoc agent, it results in poor joint performance in the presence of another ad hoc agent. A policy that explicitly accounts for the potentially unpredictable behavior of other agents and strives to cooperate with other agents would lead to better performance. In the example above, finding such a policy is as simple as introducing $\epsilon$-greedy exploration to the ad hoc agents’ action selection behavior. With that modification, both agents will continue exploring until the optimal reward of $s=1$ is achieved, in which case both agents will converge to an optimal behavior. This is because with high probability, the two ad hoc agents will learn to optimally coordinate by following a joint policy that selects one of the optimal joint actions of $(0,1)$ or $(1,0)$ at some time step, after which both agents will repeat that joint action due to the single-step agent modeling.
Figure 2: The training protocol for the proposed method, POAM. POAM adopts an independent policy gradient architecture with parameter sharing. POAM trains a single policy network, which characterizes the behavior of all controlled agents. The yellow agents are controlled by policies drawn from $U$, the set of uncontrolled agents, while the green agents are controlled by the same POAM policy. Data from both controlled and uncontrolled agents is used to train the POAM critic, $V_{\pi}$, while the actor $\pi(\theta)$ is trained on data from the controlled agents only. The actor and critic are both conditioned on a learned team encoding vector, learned via an encoder-decoder architecture.
5 Policy Optimization with Agent Modeling (POAM)
This section describes the proposed Policy Optimization with Agent Modeling (POAM) method, which trains a collection of NAHT agents that can adaptively deal with different collections of unknown teammates. POAM relies on an agent modeling network to initially build an encoding vector of the collection of teammates encountered during an interaction. Adaptive agent policies that can maximize the controlled agents’ returns are then learned by POAM by training a policy network, which receives environment observation and team encoding vector as input. Details of the training process for both agent modeling and policy networks are described in Sections 5.1 and 5.2 respectively, while an illustration of how POAM trains NAHT agents is provided in Figure 2.
5.1 Agent Modeling Network
Designing adaptive policies that enable NAHT agents to achieve optimal returns against any team of uncontrolled agents drawn from some set $U$, requires information on the encountered team’s unknown behavior. However, in the absence of prior knowledge about uncontrolled teammates’ policies, local observations from a single timestep may not contain sufficient information regarding the encountered team. To circumvent this lack of information, POAM’s agent modeling network plays a crucial role in providing team encoding vectors that characterize the behavior of teammates in the encountered team.
We identify two main criteria for desirable team encoding vectors. First, team encoding vectors should identify information regarding the unknown state and behavior of other agents in the environment. Second, team encoding vectors should ideally be computable from the sequence of local observations and actions of the team. Fulfilling both requirements provides an agent with useful information for decision-making in NAHT problems, even under partial observability.
POAM produces desirable team encoding vectors by training a model with an encoder-decoder architecture, which is illustrated by red components in Figure 2. The encoder, $f^{\text{enc}}_{\theta^{e}}:H^{i}\mapsto\mathbb{R}^{n}$, is parameterized by $\theta^{e}$ and processes the modeling agent’s history of local observations and actions up to timestep $t$, $h^{t}=\{o_{i}^{k},a_{i}^{k-1}\}_{k=1}^{t}$, to compute a team encoding vector, $e^{t}_{i}\in\mathbb{R}^{n}$. This reliance on local observations helps ensure that the agent modeling network can operate without having access to the environment state (i.e. under partial observability). We then train decoder networks, $f^{\text{dec}}_{\theta^{o}}:\mathbb{R}^{n}\mapsto O^{-i}$ and $f^{\text{dec}}_{\theta^{a}}:\mathbb{R}^{n}\mapsto\Delta(A_{-i})$, that predicts the observations and actions of the modelled agents at timestep $t$, $(o_{-i}^{t},a_{-i}^{t})$, to encourage $e^{t}_{i}$ to contain relevant information for the NAHT agents’ decision-making process. Formally, this decoder is trained to minimize the following loss function to accurately predict the modelled agents’ observations and actions:
$L_{\theta^{e},\theta^{o},\theta^{a}}(h^{t})=\Big{(}||f^{\text{dec}}_{\theta^{o}}(f^{\text{enc}}_{\theta^{e}}(h^{t}))-o_{-i}^{t}||^{2}-\text{log}(p(a^{t}_{-i};f^{\text{dec}}_{\theta^{a}}(f^{\text{enc}}_{\theta^{e}}(h^{t}))))\Big{)}.$ (2)
5.2 Policy and Value Networks
POAM relies on an actor-critic architecture to train agent policies, where the policy and networks are both conditioned on the teammate embedding produced by the encoder-decoder architecture described in Section 5.1.
The policy network, $\pi^{i}_{\theta^{p}}:H^{i}\times\mathbb{R}^{n}\mapsto\Delta(A_{i})$, uses the NAHT agent’s local observation, $o^{t}_{i}$, and the team embedding from the encoder network, $e^{t}_{i}$, to compute a policy followed by the NAHT agents. Conditioning the policy network on $e^{t}_{i}$ allows an NAHT agent to change its behaviour based on the inferred characteristics of encountered agents. When training the policy network, we also rely on a value network $V_{\theta^{c}}:H^{i}\times\mathbb{R}^{n}\mapsto\mathbb{R}$, which measures the expected returns given $h_{t}$, and $e^{t}_{i}$. The value network serves as a baseline to reduce the variance of the gradient updates and also conditions on the learned teammate embeddings, for similar reasons to the policy.
POAM then trains the policy and value networks using an approach based off the Independent PPO algorithm (Yu et al., 2022) (IPPO). IPPO is selected as the base MARL algorithm for two reasons. First, using an independent MARL method circumvents the need to deal with the changing number of agents resulting from environment openness. Second, it has been demonstrated to be effective on various MARL tasks. POAM trains the value network to produce accurate state value estimates by minimizing the following loss function:
$L_{\theta^{c}}(h^{t})=\dfrac{1}{2}\Big{(}V_{\theta^{c}}(h^{t},f^{\text{enc}}_{\theta^{e}}(h^{t}))-\hat{V}^{t}\Big{)}^{2},$ (3)
where $\hat{V}^{t}=\sum_{l=0}^{T}\gamma^{l}r^{t+l}$. The policy network is trained to minimize the PPO loss function, where the Generalized Advantage Estimator is employed (Schulman et al., 2015):
$L_{\theta^{p}}(h^{t})=\text{min}\left(r^{t}(\theta^{p})\hat{A}^{t}_{GAE},\text{clip}\left(r^{t}(\theta^{p}),1-\epsilon,1+\epsilon\right)\hat{A}^{t}_{GAE}\right)-\alpha{H}(\pi^{i}_{\theta^{p}}(.|h^{t},f^{\text{enc}}_{\theta^{e}}(h^{t}))),$ (4)
where $r^{t}(\theta^{p})=\dfrac{\pi^{i}_{\theta^{p}}(a^{t}_{i}|h^{t},f^{\text{enc}}_{\theta^{e}}(h^{t}))}{\pi^{i}_{\theta^{p_{old}}}(a^{t}_{i}|h^{t},f^{\text{enc}}_{\theta^{e}}(h^{t}))}$.
Leveraging data from uncontrolled agents
In the NAHT setting, we assume access to the joint observations and actions generated by the current team deployed in the environment, where the team consists of a mix of controlled and uncontrolled agents. POAM leverages the data from both controlled and uncontrolled agents to train the value network—in effect, treating the uncontrolled agents as exploration policies. This is a significant departure from PPO, which is a fully on-policy algorithm. However, only data from the controlled agents is used to train the policy network, as the policy update is highly sensitive to off-policy data.
6 Experiments and Results
This section presents an empirical evaluation of POAM and baseline approaches across different NAHT problems. We specifically investigate two questions. First, whether POAM leads to improved solutions for NAHT over baselines, in terms of sample efficiency or asymptoptic task return. Second, whether agent modeling enables generalization to previously unseen teammates.
We start this section by introducing the main experimental procedures and baselines in Section 6.1. Section 6.2 presents our main result that compares POAM and the baselines in NAHT problems situated across different environments. Sections 6.36.4 then provide additional analysis on the team encoding vectors produced by POAM and its ability to generalize the learned policy to previously unseen sets of teammate policies. Finally, see the Appendix for further results, hyperparameters, and implementation details. Our code will be made available upon publication.
6.1 Experimental Design
This subsection describes the experimental design, including the particular NAHT instance considered in the experiments, training procedure, experimental domain, and baselines.
6.1.1 A Practical Instantiation of NAHT
Figure 3: A practical instantiation of the NAHT problem.
Similar to AHT, the NAHT problem can be parameterized by the amount of knowledge that controlled agents have about uncontrolled agents, and whether uncontrolled agents can adapt to the behavior of controlled agents (Barrett & Stone, 2012). This section considers a practical instantiation of the NAHT problem, designed to assess whether the baseline algorithms have successfully learned to cooperate with each of the uncontrolled agents under consideration.
Recall that $U$ and $C$ denote the sets of uncontrolled and controlled agent policies, respectively, while $M$ denotes the total team size, and $N$ the number of agents sampled from $C$. For simplicity, we assume that policies in $U$ do not adapt to the controlled agents’ behavior, agents are homogeneous (so that they are interchangeable within a team), and that agents cannot communicate. While in principle, the team sampling procedure $X$ could define any probability distribution over teams constructed from $U$ and $C$, we consider an $X_{train}$ consisting of sampling $N$ uniformly from $\{1,\cdots,M-1\}$, sampling $N$ agents from $C$ and $M-N$ agents from $U$ in a uniform fashion (Figure 3). The sampling procedure described here takes place at the beginning of each episode to select a team, which is then deployed in the environment. Data generated by the deployed team (e.g. joint observations, joint actions, and rewards) are returned to the learning algorithm. See Appendix A.1.2 for further details on the evaluation procedure.
6.1.2 Generating Uncontrolled Teammates
To generate a set of uncontrolled teammates $U$, the following MARL algorithms are used to train agent teams: VDN (Sunehag et al., 2018), QMIX (Rashid et al., 2018), IQL (Tan, 1997), IPPO, and MAPPO (Yu et al., 2022). We verify that the generated team behaviors are diverse by checking that (1) teams trained by the same algorithm learn non-compatible coordination conventions, and (2) teams trained by different algorithms are not compatible (see Appendix A.2).
Let $U_{train}$ denote the set of uncontrolled teammates used to train POAM and the baseline methods. $U_{train}$ consists of five teams, where each individual team is trained via VDN, QMIX, IQL, IPPO and MAPPO, respectively. $U_{test}$ consists of a set of holdout teams trained via the same MARL algorithms, but that have not been seen during training. The experimental results reported in 6.2 and 6.3 are computed with respect to $U_{train}$ only, while the experimental results in 6.4 use $U_{test}$.
6.1.3 Experimental Domain
Experiments are conducted on the StarCraft Multi-Agent Challenge (SMAC) benchmark (Samvelyan et al., 2019). SMAC features a set of cooperative tasks, where a team of allied agents must defeat a team of enemy agents controlled by the game server. It is a partially observable domain with a continuous state space, where each agent can observe features about itself, enemy, and allied agents within some radius. The action space is discrete—each agent can choose an enemy to attack, a direction to move in, or to not perform any action. The SMAC tasks considered here are 5v6, 8v9, 10v11, 3s4z, 3s5z. The first number indicates the number of allied agents, while the second indicates the number of enemy agents; full task names and descriptions are provided in App. A.1.
6.1.4 Baselines
As NAHT is a new problem proposed by this paper, there are no prior algorithms that are directly designed for the NAHT problem. Therefore, we construct three baselines to contextualize the performance of POAM.
  • Naive MARL: various well-known MARL algorithms are considered, including both independent and centralized training with decentralized execution algorithms (Hernandez-Leal et al., 2019). The algorithms evaluated here include IQL (Claus & Boutilier, 1998), VDN (Sunehag et al., 2018), QMIX Rashid et al. (2018), IPPO, and MAPPO (Yu et al., 2022). The MARL baselines are trained in self-play and then evaluated in the NAHT setting. In the following, we report only the performance of the best naive MARL baseline.
  • Independent PPO in the NAHT setting (IPPO-NAHT): IPPO is a policy gradient MARL algorithm that directly generalizes PPO (Schulman et al., 2017) to the multi-agent setting. It was found to be surpringly effective on a variety of MARL benchmarks by Yu et al. (2022). In contrast to the naive MARL baselines, IPPO-NAHT is trained using the proposed NAHT training scheme. The variant considered here employs full parameter sharing, where the actor is trained on data only from controlled agents, but the critic is trained using data from both controlled and uncontrolled agents. The latter detail is a key algorithmic feature which POAM also employs, but is an extension from the most naive version of PPO (see Section 6.3). IPPO can be considered an ablation of POAM, where the agent modeling module is removed.
  • POAM in the AHT setting (POAM-AHT): As considered in Section 4, a natural baseline approach to the NAHT problem is to deploy AHT policies directly in the NAHT setting. To experimentally evaluate the intuition that AHT policies do not suffice for the NAHT problem setting, we consider an AHT version of POAM. POAM-AHT is trained identically to POAM, but where the number of controlled agents $N$ is always equal to 1 during training.
All methods employ full parameter sharing (Papoudakis et al., 2021b). See Appendix A.1 for details about architecture, hyperparameter tuning and algorithm implementation.
6.2 Main Results
Figure 4: POAM consistently improves over the baselines of IPPO-NAHT, POAM-AHT, and the best naive MARL baseline in all tasks, in either sample efficiency or asymptotic return. The performance of POAM-AHT and naive MARL are inconsistent across tasks. Results are computed over three trials, while the shaded region represents the 95% confidence interval.
The primary question addressed here is whether POAM improves performance on the NAHT problem setting, as compared to baselines. Figure 4 shows the learning curves of POAM and IPPO-NAHT on the SMAC tasks, and the test returns achieved by the best naive MARL baseline and POAM-AHT. See the Appendix for tables providing the performances of all naive MARL methods, across all tasks. All learning curves consist of the mean test returns across three trials, while the shaded regions reflect the 95% confidence intervals. All methods are trained for 20M time steps. We find that for all tasks, POAM outperforms baselines in either asymptotic return (5v6, 3s5z), or sample efficiency (8v9, 10v11). While POAM initially has lower sample efficiency than IPPO-NAHT, this trend is reversed after the first few million steps, after which POAM displays higher return. We attribute this to the initial costs incurred by learning team encoding vectors, which once learned, improves learning efficiency. IPPO-NAHT, which can be viewed as an ablation of POAM with no agent modeling, is the next best performing method. Although IPPO-NAHT has generally poorer sample efficiency than POAM, the method converges to the approximately the same returns on 8v9 and 10v11. While the agent modeling module of POAM provides team embedding vectors to the policy learning process, the embeddings are themselves produced from each agent’s own observation. Since no additional information is provided to POAM agents, it is unsurprising that IPPO-NAHT can converge to similar solutions as POAM, given enough training steps.
Figure 5: Evolution of the within-episode mean squared error (left) and probability of actions that were actually taken by modelled teammates (right), computed by the POAM agent over the course of training. The agent modeling performance improves over the course of an episode, as more data about teammate behavior is observed.
Finally, while the best naive MARL baseline and POAM-AHT learn good solutions on some tasks, neither method consistently performs well across all tasks. Overall, POAM discovers the most consistently performant policies compared to the other methods. We conclude that (1) the proposed learning scheme can learn effectively in the presence of multiple types of agents, and (2) agent modeling improves learning efficiency.
6.3 A Closer Look at POAM
Two key aspects of POAM are the use of a recurrent teammate modeling module, and the use of data from uncontrolled agents to train the critic. We perform further analysis of both of these aspects on the 5v6 task.
Teammate modeling performance
In the NAHT training/evaluation setting, a new set of teammates is sampled at the beginning of each episode. Therefore, an important subtask for a competent NAHT agent is to rapidly model the distribution and type of teammates at the beginning of the episode, to enable the policy to exploit that knowledge as the episode progresses. This task is especially challenging in the presence of partial observability (a property of the SMAC tasks). To address the above challenges, POAM employs a recurrent encoder, which encodes the POAM agent’s history of observations and actions to an embedding vector, and a (non-recurrent) decoder network, which predicts the egocentric observations and actions for all teammates.
A natural question then, is whether the learned teammate encoding vectors actually improve over the course of an episode. Figure 5 depicts the within-episode mean squared error (MSE) of the observation predicted by the decoder, and the within-episode probability of the action actually taken by the modelled teammates, according to the decoder’s modelled action distribution.
Each curve corresponds to a different training checkpoint. We observe the MSE rapidly converges to a low value, while the predicted probability of the taken actions has converged to 1 by time step 20. Note that the MSE is initially 0 at timestep 0 because the initial position is not randomized in SMAC. Next, we observe that the modeling performance improves over the course of training, indicated by lower prediction error at later training steps. Thus, we conclude that POAM learns accurate teammate models relatively early in the episode for the 5v6 task—successfully coping with the challenges introduced by the sampled teammates and partial observability.
Impact of data from non-controlled agents
Figure 6: Learning curves of POAM and IPPO, where the value network is trained with and without uncontrolled agents’ data (UCD).
Recall that both POAM and IPPO-NAHT update the value network using data from both the controlled and uncontrolled agents. As Figure 6 shows, this algorithmic feature results in a significant performance gain over training using on-policy data only, for both POAM and IPPO-NAHT.
6.4 Out of Distribution Generalization
Figure 7: Mean and standard error of test returns achieved by POAM and IPPO-NAHT, when paired with out-of-distribution teammates. POAM shows improved generalization to OOD teammates, as compared to IPPO-NAHT, across all tasks.
Experiments in prior sections were conducted with in-distribution teammates, i.e. teammates that the methods interacted directly with during training. This section examines whether agent modeling improves generalization to out-of-distribution (OOD) teammates, where the OOD teammates are generated by using MARL algorithms with different random seeds to generate teammates that have not been interacted with during training. [^4] We demonstrate in App. A.2 that in SMAC tasks, the same MARL algorithm with different seeds trains teams that have different coordination conventions. Hence, teams from unseen seeds display OOD behavior.
Figure 7 shows the mean and standard error of the test return achieved by POAM, compared to IPPO-NAHT, when the algorithm in question is paired with previously unseen seeds of IPPO, IQL, MAPPO, QMIX, and VDN (computed across three trials). In two out of four tasks (5v6, 3s5z), POAM has a significantly higher return than IPPO-NAHT, while the remaining two tasks exhibit a smaller improvement. For all tasks, we observe that the standard error for POAM is lower than that of IPPO-NAHT, showing that POAM performs more consistently against a range of teammates and team configurations than IPPO-NAHT.
7 Related Works
This section summarizes related work in the areas most closely related to NAHT, namely, ad hoc teamwork, zero-shot coordination, agent modeling, and cooperative MARL.
Ad Hoc Teamwork & Zero-Shot Coordination.
Prior works in ad hoc teamwork (Stone et al., 2010) and zero-shot coordination (ZSC) (Hu et al., 2020) explored methods to design adaptive agents that can optimally collaborate with unknown teammates. While they both highly resemble the NAHT problem, existing methods for AHT (Rahman et al., 2021; Mirsky et al., 2022) and ZSC (Hu et al., 2020; Lupu et al., 2021) have been limited to single-agent control scenarios. We argue that direct, naive applications of AHT and ZSC techniques to our problem of interest are ineffective—see the discussion in Section 4 and results in Section 6.
Recent research in AHT and ZSC utilizes neural networks to improve agent collaboration within various team configurations. These recent works mostly focus on two approaches. The first approach trains the agent to adapt to unknown teammates by characterizing teammates’ behavior as fixed-length vectors using neural networks and learning a policy network conditioned on these vectors (Rahman et al., 2021; Papoudakis et al., 2021a; Zintgraf et al., 2021). The second designs teammate policies that maximize the agent’s performance when collaborating with diverse teammates (Lupu et al., 2021; Rahman et al., 2021). Our work builds on the first category, extending it to control multiple agents. While this also offers a potential path for robust NAHT agents, designing teammate policies for training will be kept as future work.
Evaluating Agents’ Cooperative Capabilities.
Beyond training agents to collaborate with teammates having unknown policies, researchers have developed environments and metrics to assess cooperative abilities. The Melting Pot suite (Leibo et al., 2021; Agapiou et al., 2022) mostly evaluates controlled agents’ ability to maximize utilitarian welfare against unknown agents. However, this evaluation suite focuses on mixed-motive games where agents may have conflicting goals. This contrasts with our work’s scope of fully cooperative settings, where all agents share the same reward function.  MacAlpine & Stone (2017) also explored alternative metrics to measure agents’ cooperative capabilities while disentangling the effects of their overall skills in drop-in RoboSoccer.
Agent Modeling.
Agent modeling (Albrecht & Stone, 2018) helps us characterize other agents based on their actions. Such characterizations could attempt to infer important attributes of another agent’s decision making process such as their goals (Hanna et al., 2021) or policies (Papoudakis et al., 2021a). POAM and most existing AHT techniques rely on agent modeling to provide important teammate information for decision-making when collaborating with unknown teammates.
Cooperative MARL.
Cooperative Multi-Agent Reinforcement Learning (MARL) explores algorithms for training agent teams on fully cooperative tasks. Some existing methods focus on credit assignment and decentralized control (Foerster et al., 2018; Rashid et al., 2018). Other works in cooperative MARL also leverage parameter sharing and role assignment (e.g., Christianos et al. (2021); Yang et al. (2022)) to decide an optimal division of labor between agents. However, these techniques assume control over all existing agents during training and evaluation. This limits these techniques’ effectiveness in settings with new teammates (NAHT problem) as shown in prior work (Vezhnevets et al., 2020; Hu et al., 2020; Rahman et al., 2021).
8 Conclusion
This paper proposes and formulates the problem of $N$-agent ad-hoc teamwork (NAHT), a generalization of both AHT and MARL. It further proposes a multi-agent reinforcement learning algorithm to train NAHT agents called POAM, and develops a procedure to train and evaluate NAHT agents. POAM is a policy gradient approach that uses an encoder-decoder architecture to perform teammate modeling, and leverages data from uncontrolled agents for policy optimization. Empirical validation on StarCraft II tasks shows that POAM consistently improves over baseline NAHT approaches, in terms of sample efficiency, asymptotic return, and generalization to out-of-distribution teammates.
While this work explores teammate modeling for NAHT, other algorithmic ideas from AHT provide a rich set of future directions. These include generating teammates for the NAHT setting, alternative teammate modeling schemes, and considering communication protocols for NAHT agents. Limitations of POAM also suggest directions for future work. For instance, POAM is an independent MARL method (albeit with parameter sharing)—how might one leverage knowledge from the global state to generate improved NAHT policies? Further, POAM’s actor update is purely on-policy, and therefore cannot leverage data generated by uncontrolled agents. Future work may consider off-policy learning methods as a solution strategy.
Acknowledgments
This work has taken place in the Learning Agents Research Group (LARG) at UT Austin. LARG research is supported in part by NSF (FAIN-2019844, NRT-2125858), ONR (N00014-18-2243), ARO (E2061621), Bosch, Lockheed Martin, and UT Austin’s Good Systems grand challenge. Peter Stone serves as the Executive Director of Sony AI America and receives financial compensation for this work. The terms of this arrangement have been reviewed and approved by the University of Texas at Austin in accordance with its policy on objectivity in research.
Appendix A Appendix
A.1 Experiment Details
A.1.1 SMAC Tasks
The SMAC tasks considered in this paper are described in more detail below:
  • 5v6: stands for 5m vs 6m, five allied Marines versus six enemy Marines.
  • 8v9: stands for 8m vs 9m, eight allied Marines versus nine enemy Marines.
  • 10v11: stands for 10m vs 11m, ten allied Marines versus eleven enemy Marines.
  • 3s5z: stands for 3s vs 5z, three allied Stalkers versus five enemy Zealots.
A.1.2 Evaluation Details
The experiments in the main paper consider a particular team sampling procedure called $X_{train}$, which uniformly samples $N$ and uniformly samples agents from the sets $U$ and $C$. While the learning curves reported in the main paper are computed directly with $X_{train}$, we use an iterative and deterministic evaluation procedure for the results presented in Section 6.4 and in the Appendix. The purpose of the additional evaluation procedure presented here is to ensure that we thoroughly evaluate the trained NAHT agents for all possible values of $N$ and all possible teammates.
Given a set of uncontrolled agents $U$ and a set of agents $C$ generated by some NAHT method, we compute the following $M-N$ cross-play scores. Let $N$ be the number of agents sampled from $C$, such that $N<M$. For $N\in\{1,\cdots,M-1\}$, construct the joint policy $\bm{\pi}^{(N)}$ by selecting $N$ agents uniformly from $C$ and $M-N$ agents from $U$. Evaluate the resultant team on the task for $E$ episodes. This results in $M-1$ episode returns, on which we may compute summary statistics. If agents in $U$ and $C$ are generated by algorithm $A$ and $B$, respectively, then we call the thus computed average return, the $M-N$ cross-play score.
More concretely: each algorithm (VDN, QMIX, QMIX-NS, IQL, IQL-NS) is run $k$ times with different seeds, to generate $k$ teams of teammates to act as uncontrolled agents. For each pair of algorithms, we sample a subset of the possible seed pairs, and evaluate the teams that result from merging said seed pairs. Note that each seed actually represents a team of agents. For example, if VDN and QMIX have seeds 1 2, and 3, the cross-play evaluation might consider the seed pairings (VDN 1, QMIX 2), (VDN 2, QMIX 3), (VDN 3, QMIX 1). Given a pair of seeds (e.g. (VDN 1), (QMIX 2)), the $M-N$ cross-play score is computed as the average return generated by sweeping $N\in\{1,\cdots,M-1\}$ and evaluating the merged team that consists of selecting $N$ agents from the first team, and $M-N$ agents for $E$ episodes. In our experiments, $E=128$, and $k=3$.
Means and standard errors are computed over all seed pairings considered, as we treat each $M-N$ evaluation as an independent trial. For most experiments, three seed pairs are considered, where seeds are paired to ensure that all seeds participate in at least one evaluation. This ensures that the computational cost of the evaluation remains linear in $k$, rather than quadratic. However, note that for the OOD experiments presented in Section 6.4, we consider all possible seed pairings for the most comprehensive evaluation.
Self-play scores.
For algorithm pairs that are the same, seeds are paired with themselves. For example, the seed pairs for VDN versus VDN would be $(1,1),(2,2),(3,3)$. The reason for this is that we are interested in the performance of agents when paired with known teammates. This is the model self-play score, in contrast to the algorithm self-play score (Albrecht et al., 2024).
A.1.3 Algorithm Implementation
The experiments in this paper use algorithm implementations from the ePyMARL codebase (Papoudakis et al., 2021b). The value-based methods are used without modification (i.e. IQL, VDN, QMIX), but we implement our own version of policy gradient methods (IPPO, MAPPO), based on the implementation of (Yu et al., 2022).
All methods employ recurrent actors and critics, with full parameter sharing, i.e. all agents are controlled by the same policy, where the agent id is input to the policy/critic networks, to allow behavioral differentiation between agents. For policy gradient methods, the policy architecture is two fully connected layers, followed by an RNN (GRU) layer, followed by an output layer. Each layer has 64 neurons with ReLU activation units, and employs layer normalization. The critic architecture is the same as the policy architecture. The value-based networks employ the same architecture, except that there is only a single fully connected layer before the RNN layers, and layer normalization is not used (following the ePyMARL implementation). Please consult the code base for full implementation details.
Hyperparameters.
For the value-based methods, default hyperparameters are used. We tune the hyperparameters of the policy gradient methods on the 5v6 task, and apply those parameters directly to the remaining SMAC tasks. The hyperparameters considered for policy gradient algorithms are given in Table 1. POAM adopts the same hyperparameters as IPPO where applicable. We also tuned additional hyperparameters specific to POAM (see Table 2).
Algorithm Buffer size Epochs Minibatches Entropy Clip Clip value loss
IPPO 128, 256, 512 1, 4, 10 1, 3 0.01, 0.03, 0.05 0.01, 0.05, 0.1 no, yes
MAPPO 64, 128, 256 4, 10 1, 3 0.01, 0.03, 0.05, 0.07 0.05, 0.1, 0.2 no, yes
Table 1: Hyperparameters evaluated for the policy gradient algorithms. Selected values are bolded
Algorithm ED epochs ED Minibatches ED LR
POAM 1, 5, 10 1, 2 0.0005, 0.005
Table 2: Additional hyperparameters evaluated for POAM; note that ED stands for encoder-decoder. Selected values are bolded.
A.2 Supplemental Results
A.2.1 On coordination conventions in StarCraft II
The experimental procedure detailed in Section 6.1 generates diverse teammates in SMAC by using MARL algorithms to train multiple teams of agents. Two underlying assumptions of the procedure are that (1) teams trained by the same algorithm learn non-compatible coordination conventions, and (2) teams trained by different algorithms are not compatible. Both points are experimentally verified in this section.
Self-play with MARL algorithms.
For the SMAC tasks under consideration, teams trained using the same MARL algorithm, but different seeds, can converge to distinct coordination conventions. Figure 8 demonstrates this by depicting the return of teams trained together (matched seeds) versus those of teams that were not trained together (mismatched seeds), across all naive MARL algorithms (IPPO, IQL, MAPPO, QMIX, VDN) and tasks considered. Overall, the returns of the teams that were trained together are higher than those not trained together.
The general phenomenon has been previously observed and exploited by prior works in ad hoc teamwork (Strouse et al., 2022). This paper takes advantage of this to generate a set of diverse teammates for the StarCraft II experiments. We select tasks where the effect is significant, to ensure that there are distinct coordination behaviors for POAM to model. Tasks that were considered but subsequently ruled out for this reason include 3m vs 3m, 8m vs 8m, and 6h vs 8z.
Figure 8: Agent teams that were trained together using the same algorithm (matched seeds) have higher returns than teams that were not trained together but were trained with the same algorithm (mismatched seeds).
Cross-play with MARL algorithms
Tables 9, 10, 11, and 12 display the full cross-play results for all MARL algorithms on all tasks. Note that the tables are reflected across the diagonal axis for viewing ease. The values on the off-diagonal (i.e., where the two algorithms are not the same) are the cross-play score, while the values shown on the diagonal are self-play scores, computed as described in App. A.1.2. The cross-play and self-play scores displayed are means and standard errors. The rightmost column reflects the row average and standard deviation of the cross-play (xp) scores corresponding to the test set of VDN, QMIX, IQL, IPPO, and MAPPO, excluding the self-play score. Thus, the rightmost column reflects how well on average the row algorithm can generalize to unfamiliar teams.
Overall, we find that MARL algorithms perform significantly better in self-play than cross-play. We note that there are a few exceptions (e.g., IPPO vs QMIX on 8m), and that the cross-play and self-play scores are much closer on 10v11, but overall the trend is consistent across tasks.
Figure 9: Cross-play table for the 5v6 task.
Figure 10: Cross-play table for the 8m task.
Figure 11: Cross-play table for the 3s5z task.
Figure 12: Cross-play table for the 10v11 task.
A.3 Cross-play performance as a function of the number of agents.
To provide more insight on how a non-NAHT team performance changes as the number of unseen agents increases, Figure 13 plots QMIX and IPPO’s test scores on 5v6 when paired with various methods, as a function of the number of unseen agents. Note that $k=0$ corresponds to teams entirely composed of the evaluated agents, while $k=5$ corresponds to teams entirely composed of the unseen teammates, and are actually self-play scores, whereas the other points are cross-play scores. These points appear on the plot to contextualize the performance of the mixed teams.
QMIX and IPPO have similar returns in self-play, but comparing the two figures, we observe a difference in the rate at which the cross-play returns decline. In particular, QMIX’s performance declines most around $k=3$ across a variety of methods, while IPPO’s performance is minimized at $k=2$.
Figure 13: Mean and std. error of QMIX (left) and IPPO (right) test returns on 5v6, in the presence of uncontrolled agents generated by various MARL learning algorithms, as the number of unseen agents increases from $k=0$ to a maximum of $k=5$. The unseen agents are generated by the second method in the legend name.
A.4 Further Discussion of the Motivating Example
This section considers an extension of the motivating example discussed in Section 4; namely, whether considering longer agent modeling windows is sufficient to allow the ad hoc agents to solve the $s=1$ game without coordination.
Consider a modification of the ad hoc agents’ policies such that instead of picking the best response to teammate actions in the last step, they consider the last $t$ actions of each teammate, and pick the best response with respect to the expected actions. In this extended scenario, assume there are $M-1$ (uncontrolled) agents which always pick 00. Assume that the $M-1$ always-zero agents play without the ad hoc agent for $t$ steps. Then, the ad hoc agent $a_{h}$ is introduced into the game in the $t+1$-th step. In this scenario, $a_{h}$ will learn that all the other agents always pick zero, and will pick the optimal action $1$ accordingly.
Now consider the setting where there are $M-2$ always-zero agents playing without the ad hoc agent for $t$ steps, and two ad hoc agents $a_{h}^{1}$ and $a_{h}^{2}$ are introduced into the game at the $t+1$-th step. Following the same modeling policy, the two ad hoc agents $a_{h}^{1}$ and $a_{h}^{2}$ will once again both learn that their teammates follow an always-zero policy, and therefore learn to pick the $1$ action — a suboptimal decision, as $s=2$.
In this setting, even if $a_{h}^{1}$ and $a_{h}^{2}$ modeled teammate behavior on a sliding window of length $t$ rather than every $t$ steps in batch form, they will not converge to optimal behavior. To see why, consider the actions of $a_{h}^{1}$ and $a_{h}^{2}$ in an arbitrary window of size $w$. Since they begin in unison, are deterministic, and consider the same window size, their actions will be identical for all the $t$ steps of this window. For the first $t$ steps, lacking history, both agents will pick action $1$, as we have shown above. After $k$ steps, they will both switch to 00, as we have also shown. Now, however, if they are both modeling agent action distributions using a rolling window of the same size $t$, after $t/2+1$ additional steps, they will reach a point in which both agents have chosen 00 more than 50% of the time in the last $t$ timesteps.
At that point, both agents will once again simultaneously switch their actions to $1$ to maximize the likelihood of winning the game. As a result $S$ would once again flip from 00 to $2$. Using a longer window for modeling action does not mitigate the problem of inadequately modeling other ad hoc agents, which leads to an inability to cooperate.
A.5 Computing Infrastructure
All value-based algorithms (e.g. QMIX, VDN, IQL) were run without parallelized training, while policy gradient algorithms were run with parallelized training. The servers used in our experiments ran Ubuntu 18.04 with the following configurations:
  • Intel Xeon CPU E5-2630 v4; Nvidia Titan V GPU.
  • Intel Xeon CPU E5-2698 v4; Nvidia Tesla V100-SXM2 GPU.
  • Intel Xeon Gold 6342 CPU; Nvidia A40 GPU.
  • Intel Xeon Gold 6342 CPU; Nvidia A100 Gpu.

### v2

@@
 Abstract
-Current approaches to learning cooperative behaviors in multi-agent settings assume relatively restrictive settings.
+Current approaches to learning cooperative multi-agent behaviors assume relatively restrictive settings.
 In standard fully cooperative multi-agent reinforcement learning, the learning algorithm controls all agents in the scenario, while in ad hoc teamwork, the learning algorithm usually assumes control over only a single agent in the scenario.
@@
 For example, in an autonomous driving scenario, a company might train its cars with the same learning algorithm, yet once on the road, these cars must cooperate with cars from another company.
-Towards generalizing the class of scenarios that cooperative learning methods can address, we introduce $N$-agent ad hoc teamwork, in which a set of autonomous agents must interact and cooperate with dynamically varying numbers and types of teammates at evaluation time.
+Towards expanding the class of scenarios that cooperative learning methods may optimally address, we introduce $N$-agent ad hoc teamwork (NAHT), where a set of autonomous agents must interact and cooperate with dynamically varying numbers and types of teammates.
 This paper formalizes the problem, and proposes the Policy Optimization with Agent Modelling (POAM) algorithm.
 POAM is a policy gradient, multi-agent reinforcement learning approach to the NAHT problem, that enables adaptation to diverse teammate behaviors by learning representations of teammate behaviors.
-Empirical evaluation on StarCraft II tasks shows that POAM improves cooperative task returns compared to baseline approaches, and enables out-of-distribution generalization to unseen teammates.
+Empirical evaluation on tasks from the multi-agent particle environment and StarCraft II shows that POAM improves cooperative task returns compared to baseline approaches, and enables out-of-distribution generalization to unseen teammates.
 1 Introduction
-Cooperative multi-agent reinforcement learning (MARL) is a paradigm for learning agent teams that solve a common task via interaction with each other and the environment (Littman, 1994).
+Advances in multi-agent reinforcement learning (MARL) [3] have enabled agents to learn solutions to various problems in zero-sum games, social dilemmas, adversarial team games, and cooperative tasks [35, 8, 18, 20, 32].
+Within MARL, cooperative multi-agent reinforcement learning (CMARL) is a paradigm for learning agent teams that solve a common task via interaction with each other and the environment [21, 27, 44].
+Recent CMARL methods have been able to learn impressive examples of cooperative behavior from scratch in controlled settings, where all agents are controlled by the same learning algorithm [32, 36, 4].
 A related paradigm for learning cooperative behavior is ad hoc teamwork (AHT).
-In contrast to MARL, the objective of AHT is to create a single agent policy that can collaborate with unknown teammates to solve a common task (Mirsky et al., 2022; Stone et al., 2010).
-While impressive, these advances have largely been confined to scenarios in which either complete control over all agents is assumed, or only a single agent is adapted for cooperation (Rashid et al., 2018; Son et al., 2019; Lin et al., 2023; Baker et al., 2019).
+In contrast to CMARL, the objective of AHT is to create a single agent policy that can collaborate with previously unknown teammates to solve a common task [26, 37].
+While a large and impressive body of work on CMARL and AHT exists, the current literature has largely examined scenarios in which either complete control over all agents is assumed, or only a single agent is adapted for cooperation [32, 36, 20, 4, 14].
+Even learning methods for handling cooperative tasks in open multiagent systems [44], which encompass one of the most challenging settings where agents may enter or leave the system anytime, either operate assuming full control over all agents [16, 22] or only a single adaptive agent [30, 17, 31].
 However, real-world collaborative scenarios—e.g.
 search-and-rescue, or robot fleets for warehouses—might demand agent subteams that are able to collaborate with unfamiliar teammates that follow different coordination conventions.
-Towards producing agent teams that are more flexible and applicable to realistic cooperative scenarios, this paper proposes the novel problem setting of $N$-agent ad hoc teamwork (NAHT), in which a set of autonomous agents must interact with an uncontrolled set of teammates to perform a cooperative task.
+Towards producing agent teams that are more flexible and applicable to realistic cooperative scenarios, this paper formalizes the problem setting of $N$-agent ad hoc teamwork (NAHT), in which a set of autonomous agents must interact with an uncontrolled set of teammates to perform a cooperative task.
 When there is only a single ad hoc agent, NAHT is equivalent to AHT.
-On the other hand, when all ad hoc agents were produced by some MARL algorithm and there are no uncontrolled teammates, NAHT is equivalent to MARL.
-Thus, the proposed problem setting generalizes both MARL and AHT.
-Drawing from ideas in both MARL and AHT, we introduce Policy Optimization with Agent Modelling (POAM).
+On the other hand, when all ad hoc agents are jointly trained by the same algorithm and there are no uncontrolled teammates, NAHT is equivalent to CMARL.
+Thus, the proposed problem setting generalizes both CMARL and AHT.
+Drawing from ideas in both CMARL and AHT, we introduce Policy Optimization with Agent Modelling (POAM).
 POAM is a policy-gradient based approach for learning cooperative multi-agent team behaviors, in the presence of varying numbers and types of teammate behaviors.
 It consists of (1) an agent modeling network that generates a vector characterizing teammate behaviors, and (2) an independent actor-critic architecture, which conditions on the learned teammate vectors to enable adaptation to a variety of potential teammate behaviors.
-Empirical evaluation on StarCraft II tasks shows that POAM learns to coordinate with a changing number of teammates of various types, with higher competency than MARL, AHT, and NAHT baseline approaches.
-An evaluation with out-of-distribution teammates also reveals that POAM’s agent modeling module enables improved generalization to out-of-distribution teammates, compared to baseline without agent modeling.
-To summarize, this paper makes the following contributions:
-  • It proposes, motivates, and formalizes the novel problem setting of N-agent ad hoc teamwork.
-  • It introduces Policy Optimization with Agent Modelling (POAM), an algorithm that leverages ideas from AHT and MARL to enable learning cooperative team behaviors whilst dealing with diverse behaviors from uncontrolled teammates.
-  • It empirically demonstrates that POAM enables improved cooperative task returns over baselines, and generalization to out-of-distribution teammates.
-Figure 1: Left: MARL algorithms assume full control over all $M$ agents in a cooperative scenario.
-Center: AHT algorithms often assume that only a single agent is controlled by the learning algorithm, while the other $M-1$ agents are uncontrolled and can have a diverse, unknown set of behaviors.
-Right: NAHT, the paradigm proposed by this paper, assumes that $N$ agents (where $N$ can vary) are controlled by the learning algorithm, while the remaining $M-N$ agents are uncontrolled.
-2 Background
+Empirical evaluation on multi-agent particle environment (MPE) and StarCraft II tasks shows that POAM learns to coordinate with a changing number of teammates of various types, with higher competency than CMARL, AHT, and naive NAHT baseline approaches.
+An evaluation with out-of-distribution teammates also reveals that POAM’s agent modeling module improves generalization to out-of-distribution teammates, compared to baselines without agent modeling.
+Figure 1: Left: CMARL algorithms assume full control over all $M$ agents in a cooperative scenario.
+Center: AHT algorithms assume that only a single agent is controlled by the learning algorithm, while the other $M-1$ agents are uncontrolled and can have a diverse, unknown set of behaviors.
+Right: NAHT, the paradigm proposed by this paper, assumes that a potentially varying $N$ agents are controlled by the learning algorithm, while the remaining $M-N$ agents are uncontrolled.
+2 Background and Notation
 […]
 3 NAHT Problem Formulation
-Drawing from the goals of MARL and AHT (Stone et al., 2010), the goal of $N$-agent ad-hoc teamwork is to create a set of autonomous agents that are able to efficiently collaborate with both known and unknown teammates to maximize return on a task.
+Drawing from the goals of MARL and AHT [37], the goal of $N$-agent ad hoc teamwork is to create a set of autonomous agents that are able to efficiently collaborate with both known and unknown teammates to maximize return on a task.
 The goal is formalized below.
-Let $C$ denote a set of ad-hoc agents.
+Let $C$ denote a set of ad hoc agents.
 If the policies of the agents in $C$ are generated by an algorithm, we say that the algorithm controls agents in $C$.
@@
 Let $U$ denote a set of uncontrolled agents, which we define as all agents in the environment not included in $C$ [^2] The original AHT problem statement used the terms “known” and “unknown” agents.
-However, it is common for modern AHT learning methods to assume some knowledge about the “unknown” teammates during training (Mirsky et al., 2022).
+However, it is common for modern AHT learning methods to assume some knowledge about the “unknown” teammates during training [26].
 Thus, we instead employ the terms controlled and uncontrolled.
-We adopt an analogous assumption to Stone et al.
-(2010)—namely, that agents in $U$ satisfy a minimal level of competency in a team setting.
+Following Stone et al., we assume that agents in $U$ are not adversarially minimizing the objective of agents in $C$.
 We model an open system of interaction, in which a random selection of $M$ agents from sets $C$ and $U$ must coordinate to perform task $r$.
 For illustration, consider a warehouse staffed by robots developed by companies $A$ and $B$, where there is a box-lifting task that requires three robots to accomplish.
-If Company A’s robots are ad-hoc agents (corresponding to $C$), then some robots from Company A could collaborate with robots from Company B (corresponding to $U$) to accomplish the task, rather than requiring that all three robots come exclusively from $A$ or $B$.
-Motivated thus, we introduce a team sampling procedure $X$.
+If Company A’s robots are controlled agents (corresponding to $C$), then some robots from Company A could collaborate with robots from Company B (corresponding to $U$) to accomplish the task, rather than requiring that all three robots come exclusively from $A$ or $B$.
+Motivated thus, we introduce a team sampling procedure $X(U,C)$.
 At the beginning of each episode, $X$ samples a team of $M$ agents by first sampling $N<M$, then sampling $N$ agents from the set $C$ and $M-N$ agents from $U$.
-We define $X$ more formally below.
+We restrict consideration to teams containing at least one controlled agent, i.e $N\geq 1$.
+We consider $X$ a problem parameter that is not under the control of any algorithm for generating ad hoc teammates, analogous to the transition function of the underlying Dec-POMDP.
+A more explicit definition of $X$ is provided in Appendix A.1.
+Without loss of generality, let $C(\theta)=\{\pi^{i}_{\theta}(.|s)\}_{i=1}^{N}$ denote a set of $N$ controlled agent policies parameterized by $\theta$, such that a learning algorithm might optimize for $\theta$.
+Let the $\bm{\pi}^{(M)}$ indicate a team of $M$ agents and $\bm{\pi}^{(M)}\sim X(U,C)$ indicate sampling such a team from $U$ and $C$ via the team sampling procedure.
+The objective of the NAHT problem is to find parameters $\theta$, such that $C(\theta)$ maximizes the expected return in the presence of teammates from $U$:
+$\max_{\theta}\left(\mathbb{E}_{\bm{\pi}^{(M)}\sim X(U,C(\theta))}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}\right]\right).$ (1)
+Challenges of the NAHT problem include: (1) coordinating with potentially unknown teammates (generalization), and (2) coping with a varying number of uncontrolled teammates (openness).
+4 The Need for Dedicated NAHT Algorithms
+Having introduced the NAHT problem, a natural question to consider is whether AHT solutions may optimally address NAHT problems.
+If so, then there would be little need to consider the NAHT problem setting.
+For instance, a simple yet reasonable approach consists of directly using an AHT policy to control as many agents as required in an NAHT scenario.
+This section illustrate the limitations of the aforementioned approach by giving a concrete example of a matrix game where (1) an AHT policy that is learned in the AHT ($N=1$) scenario is unlikely to do well in an NAHT scenario where $N=2$, and (2) even an optimal AHT policy is suboptimal in the $N=2$ setting.
+Define the following simple game for $M$ agents: at each turn, each agent $a_{i}$ picks one bit $b_{i}\in\{0,1\}$; at the end of each turn, all the bits are summed $s=\sum_{i}b_{i}$.
+The team wins if the sum of the chosen bits is exactly $1$.
+We denote the probability of winning by $P(s=1)$.
+Suppose the uncontrolled agents follow a policy that independently selects $1$ with probability $p=\frac{1}{M}$.
+[^3] Lemma A.1 shows that this is the optimal value of $p$ for a team composed of such agents.
+In the following, we consider the three agent case, $M=3$, for simplicity.
+In the AHT problem setting, a learning algorithm assumes control of only a single agent.
+Let $p_{\text{aht}}$ denote the probability with which the AHT agent selects 1.
+Given the aforementioned team of uncontrolled agents, we show that any value of $p_{\text{aht}}$ results in the same $P(s=1)$, i.e.
+all possible policies for the AHT agent are optimal.
+This occurs because the probability of winning, $P(s=1)=\frac{4}{9}$, is independent of $p_{\text{aht}}$ (Lemma A.2).
+Next, consider an NAHT scenario where a learning algorithm must define the actions of two out of three agents.
+Suppose that the same AHT policy is used to control both agents: both agents select 1 with probability $p_{\text{aht}}$.
+Above, we demonstrated that an AHT algorithm trained in the $N=1$ scenario could result in learning any $p_{\text{aht}}$.
+However, in the $N=2$ setting, we show that the optimal AHT policy $p_{\text{aht}}=\frac{1}{3}$ and the winning probability for this policy is $P(s=1)=\frac{4}{9}$ (Lemma A.3).
+Finally, we show that there exists an NAHT policy that controls both agents and obtains a higher winning probability.
+Consider the policy where one controlled agent always plays 0, while the other plays 1 with probability $p_{\text{naht}}$.
+We show that the optimal $p_{\text{naht}}=1$, and the probability of winning $P(s=1)=\frac{2}{3}>\frac{4}{9}$ (Lemma A.4).
+Thus, we have exhibited an NAHT scenario where an AHT policy that is optimal when $N=1$, performs worse than a simple NAHT joint policy in the $N=2$ setting.
+5 Policy Optimization with Agent Modeling (POAM)
+Figure 2: POAM trains a single policy network, which characterizes the behavior of all controlled agents (green), while uncontrolled agents (yellow) are drawn from $U$.
+Data from both controlled and uncontrolled agents is used to train the critic, $V_{\pi}$ while the actor $\pi(\theta)$ is trained on data from the controlled agents only.
+The actor and critic are both conditioned on a learned team encoding vector.
+This section describes the proposed Policy Optimization with Agent Modeling (POAM) method to trains a collection of NAHT agents that can adaptively deal with different collections of unknown teammates.
+POAM relies on an agent modeling network to initially build an encoding vector of the collection of teammates encountered during an interaction.
+Adaptive agent policies that can maximize the controlled agents’ returns are then learned by training a policy conditioned on the environment observation and team encoding vector.
+The training processes for agent modeling and policy networks are described in Sections 5.1 and 5.2 respectively, while an illustration of how POAM trains NAHT agents is provided in Figure 2.
+5.1 Agent Modeling Network
+Designing adaptive policies that enable NAHT agents to achieve optimal returns against any team of uncontrolled agents drawn from some set $U$, requires information on the encountered team’s unknown behavior.
+However, in the absence of prior knowledge about uncontrolled teammates’ policies, local observations from a single timestep may not contain sufficient information regarding the encountered team.
+To circumvent this lack of information, POAM’s agent modeling network plays a crucial role in providing team encoding vectors that characterize the observed behavior of teammates in the encountered team.
+We identify two main criteria for desirable team encoding vectors.
+First, team encoding vectors should identify information regarding the unknown state and behavior of other agents in the environment (both controlled and uncontrolled).
+Second, team encoding vectors should ideally be computable from the sequence of local observations and actions of the team.
+Fulfilling both requirements provides an agent with useful information for decision-making in NAHT problems, even under partial observability.
+POAM produces desirable team encoding vectors by training a model with an encoder-decoder architecture, which is illustrated by red components in Figure 2.
+The encoder, $f^{\text{enc}}_{\theta^{e}}:H^{i}\mapsto\mathbb{R}^{n}$, is parameterized by $\theta^{e}$ and processes the modeling agent’s history of local observations and actions up to timestep $t$, $h^{t}=\{o_{i}^{k},a_{i}^{k-1}\}_{k=1}^{t}$, to compute a team encoding vector, $e^{t}_{i}\in\mathbb{R}^{n}$.
+This reliance on local observations helps ensure that the agent modeling network can operate without having access to the environment state that is unavailable under partial observability.
+The decoder networks, $f^{\text{dec}}_{\theta^{o}}:\mathbb{R}^{n}\mapsto O^{-i}$ and $f^{\text{dec}}_{\theta^{a}}:\mathbb{R}^{n}\mapsto\Delta(A_{-i})$, are trained to predict the observations and actions of the modeled agents at timestep $t$, $(o_{-i}^{t},a_{-i}^{t})$, to encourage $e^{t}_{i}$ to contain relevant information for the NAHT agents’ decision-making process.
+Formally, this decoder is trained to minimize the following loss function to accurately predict the modeled agents’ observations and actions:
+$L_{\theta^{e},\theta^{o},\theta^{a}}(h^{t})=\Big{(}||f^{\text{dec}}_{\theta^{o}}(f^{\text{enc}}_{\theta^{e}}(h^{t}))-o_{-i}^{t}||^{2}-\text{log}(p(a^{t}_{-i};f^{\text{dec}}_{\theta^{a}}(f^{\text{enc}}_{\theta^{e}}(h^{t}))))\Big{)}.$ (2)
+5.2 Policy and Value Networks
+POAM relies on an actor-critic approach to train agent policies, where the policy and critic are both conditioned on the teammate embedding described in Section 5.1.
+The policy network, $\pi^{i}_{\theta^{p}}:H^{i}\times\mathbb{R}^{n}\mapsto\Delta(A_{i})$, uses the NAHT agent’s local observation, $o^{t}_{i}$, and the team embedding from the encoder network, $e^{t}_{i}$, to compute a policy followed by the NAHT agents.
+Conditioning the policy network on $e^{t}_{i}$ allows an NAHT agent to change its behaviour based on the inferred characteristics of encountered agents.
+When training the policy network, we also rely on a value network $V_{\theta^{c}}:H^{i}\times\mathbb{R}^{n}\mapsto\mathbb{R}$, which measures the expected returns given $h_{t}$, and $e^{t}_{i}$.
+The value network serves as a baseline to reduce the variance of the gradient updates and also conditions on the learned teammate embeddings, for similar reasons to the policy.
+POAM then trains the policy and value networks using an approach based off the Independent PPO algorithm [43] (IPPO).
+IPPO is selected as the base MARL algorithm for two reasons.
+First, using an independent MARL method circumvents the need to deal with the changing number of agents resulting from environment openness.
+Second, it has been demonstrated to be effective on various MARL tasks.
+POAM trains the value network to produce accurate state value estimates by minimizing the following loss function:
+$L_{\theta^{c}}(h^{t})=\dfrac{1}{2}\Big{(}V_{\theta^{c}}(h^{t},f^{\text{enc}}_{\theta^{e}}(h^{t}))-\hat{V}^{t}\Big{)}^{2},$ (3)
+where $\hat{V}^{t}=\sum_{l=0}^{T}\gamma^{l}r^{t+l}$. The policy network is trained to minimize the PPO loss function [34].
+Leveraging data from uncontrolled agents
+In the NAHT setting, we assume access to the joint observations and actions generated by the current team deployed in the environment, where the team consists of a mix of controlled and uncontrolled agents.
+POAM leverages data from both controlled and uncontrolled agents to train the value network—in effect, treating the uncontrolled agents as exploration policies.
+This aspect of POAM is a significant departure from PPO, which is a fully on-policy algorithm.
+However, only data from the controlled agents is used to train the policy network, as the policy update is highly sensitive to off-policy data.
+6 Experiments and Results
+This section presents an empirical evaluation of POAM and baseline approaches across different NAHT problems. We investigate three questions and foreshadow the conclusions:
+  • Q1: Does POAM lead to better sample efficiency and/or asymptotic return than baselines? (Usually)
+  • Q2: Can we verify that the two key ideas of POAM—agent modelling and use of data from uncontrolled agents—work as desired and contribute positively towards POAM’s performance? (Yes)
+  • Q3: Does agent modelling improve generalization to previously unseen teammates? (Yes)
+Full implementation details, including all hyperparameter values, along with additional empirical results, appear in Appendix.
+Our code is included with the submission and will be open-sourced upon publication.
+6.1 Experimental Design
+This subsection describes the experimental design, including the particular NAHT problem setting, training procedure, experimental domain, and baselines.
+A Practical Instantiation of NAHT
+Similarly to AHT, the NAHT problem can be parameterized by the amount of knowledge that controlled agents have about uncontrolled agents, and whether uncontrolled agents can adapt to the behavior of controlled agents [5].
+Furthermore, as a direct result of the fact that an NAHT algorithm may control more than a single agent, the NAHT problem may also be parameterized by whether the controlled agents are homogeneous, whether they can communicate, and what the team sampling procedure is.
+While the fully general problem setting allows for heterogeneous, communicating, controlled agents that have no knowledge of the uncontrolled agents, as a first step, this paper focuses on a special case of the NAHT problem, where agents are homogeneous, non-communicating, may learn about uncontrolled agents via interaction, and where the team sampling procedure consists of a uniform random sampling scheme from $U$ and $C$ (see Appendix A.3.1 for details).
+This case is designed primarily to assess whether controlled subteams may outperform independent controlled agents, when cooperating with multiple types of uncontrolled agents.
+We leave consideration of broader NAHT scenarios for future work.
+Generating Uncontrolled Teammates
+To generate a set of uncontrolled teammates $U$, the following MARL algorithms are used to train agent teams: VDN [39], QMIX [32], IQL [40], IPPO, and MAPPO [43].
+We verify that the generated team behaviors are diverse by checking that (1) teams trained by the same algorithm learn non-compatible coordination conventions, and (2) teams trained by different algorithms are not compatible (Appendix A.4).
+This emergence of diverse teammate behaviors from training agents using different MARL algorithms under different seeds aligns with the findings from Strouse et al.
+[38].
+Let $U_{train}$ denote the set of uncontrolled teammates used to train POAM and the baseline methods.
+$U_{train}$ consists of five teams, where each individual team is trained via VDN, QMIX, IQL, IPPO and MAPPO, respectively.
+$U_{test}$ consists of a set of holdout teams trained via the same MARL algorithms, but that have not been seen during training.
+The experimental results reported in 6.2 and 6.3 are computed with respect to $U_{train}$ only, while the experimental results in 6.4 use $U_{test}$.
+Experimental Domains
+Experiments are conducted on a predator-prey mpe-pp task implemented within the multi-agent particle environment  [23], and the 5v6, 8v9, 10v11, 3s5z tasks from the StarCraft Multi-Agent Challenge (SMAC) benchmark [33].
+On the mpe-pp task, three predators must cooperatively pursue a pretrained prey agent.
+The team receives a reward of +1 per time step that two or more predators collide with the prey.
+On the SMAC tasks, a team of allied agents must defeat a team of enemy agents controlled by the game server.
+For each task, the first number in the task name indicates the number of allied agents, while the second indicates the number of enemy agents.
+The team is rewarded for defeating enemies, with a large bonus for defeating all enemies.
+See Appendix A.3.2 for full details.
+Baselines
+As NAHT is a new problem proposed by this paper, there are no prior algorithms that are directly designed for the NAHT problem.
+Therefore, we construct three baselines to contextualize the performance of POAM.
+All methods employ full parameter sharing [29].
+  • Naive MARL: various well-known MARL algorithms are considered, including both independent and centralized training with decentralized execution algorithms [13].
+The algorithms evaluated here include IQL [10], VDN [39], QMIX [32], IPPO, and MAPPO [43].
+The MARL baselines are trained in self-play and then evaluated in the NAHT setting.
+In the following, only the performance of the best naive MARL baseline is reported.
+  • Independent PPO in the NAHT setting (IPPO-NAHT): IPPO is a policy gradient MARL algorithm that directly generalizes PPO [34] to the multi-agent setting.
+It was found to be surprisingly effective on a variety of MARL benchmarks [43].
+In contrast to the naive MARL baselines, IPPO-NAHT is trained using the NAHT training scheme presented in Section 6.1.
+The variant considered here employs full parameter sharing, where the actor is trained on data only from controlled agents, but the critic is trained using data from both controlled and uncontrolled agents.
+The latter detail is a key algorithmic feature which POAM also employs, but is an extension from the most naive version of PPO (see Section 6.3).
+IPPO can be considered an ablation of POAM, where the agent modeling module is removed.
+  • POAM in the AHT setting (POAM-AHT): As considered in Section 4, a natural baseline approach to the NAHT problem is use AHT algorithms that train only a single controlled agent, and copy these policies as many times as needed in the NAHT setting.
+To evaluate the intuition that AHT policies do not suffice for the NAHT problem setting, we consider an AHT version of POAM that is trained identically to POAM, but where the number of controlled agents $N=1$ during training.
+6.2 Main Results
+Figure 3: POAM consistently improves over the baselines of IPPO-NAHT, POAM-AHT, and the best naive MARL baseline in all tasks, in either sample efficiency or asymptotic return.
+Results are computed over three trials, while the shaded region represents the 95% confidence interval.
+This section addresses Q1—that is, whether POAM improves asymptotic return and/or sample efficiency, compared to baselines.
+Figure 3 shows the learning curves of POAM and IPPO-NAHT, and the test returns achieved by the best naive MARL baseline and POAM-AHT, on all tasks.
+[^4] See the Appendix for tables providing the performances of all naive MARL methods, across all tasks.
+All learning curves consist of the mean test returns across three trials, while the shaded regions reflect the 95% confidence intervals.
+We find that for all tasks, POAM outperforms baselines in terms of asymptotic return for three out of five tasks (pp, 5v6, 3s5z).
+For all tasks, POAM’s initial sample efficiency is similar to that of IPPO-NAHT for the first few million steps of training, after which POAM displays higher return.
+We attribute this to the initial costs incurred by learning team encoding vectors, which once learned, improves learning efficiency.
+IPPO-NAHT, which can be viewed as an ablation of POAM with no agent modeling, is the next best performing method.
+Although IPPO-NAHT has generally poorer sample efficiency than POAM, the method converges to approximately the same returns on two out of five tasks (8v9 and 10v11).
+While the agent modeling module of POAM provides team embedding vectors to the policy learning process, the embeddings are themselves produced from each agent’s own observation.
+Since no additional information is provided to POAM agents, it is unsurprising that IPPO-NAHT can converge to similar solutions as POAM, given enough training steps.
+Figure 4: Evolution of the within-episode mean squared error (left) and the within-episode probability of actions of modelled teammates (right), computed by the POAM agent over the course of training on the mpe-pp task.
+Finally, while the best naive MARL baseline and POAM-AHT learn good solutions on some tasks, neither method consistently performs well across all tasks.
+Overall, POAM discovers the most consistently performant policies compared to baseline methods, in a sample-efficient manner.
+We conclude that (1) agent modeling improves learning efficiency, and (2) direct duplication of AHT-trained policies is less effective than methods that co-train agents for NAHT.
+6.3 A Closer Look at POAM
+Two key aspects of POAM are the use of a recurrent teammate modeling module, and the use of data from uncontrolled agents to train the critic (Q2).
+We perform further analysis of both aspects on the mpe-pp and 5v6 tasks; results on 5v6 may be found in Appendix A.4.
+Teammate modeling performance
+In the NAHT training/evaluation setting, a new set of teammates is sampled at the beginning of each episode.
+Therefore, an important subtask for a competent NAHT agent is to rapidly model the distribution and type of teammates at the beginning of the episode, to enable the policy to exploit that knowledge as the episode progresses.
+This task is especially challenging in the presence of partial observability (a property of the SMAC tasks).
+To address the above challenges, POAM employs a recurrent encoder, which encodes the POAM agent’s history of observations and actions to an embedding vector, and a (non-recurrent) decoder network, which predicts the egocentric observations and actions for all teammates.
+A natural question then, is whether the learned teammate encoding vectors actually improve over the course of an episode.
+Figures 4 and 9 depict the within-episode mean squared error (MSE) of the observation predicted by the decoder, and the within-episode probability of the action actually taken by the modelled teammates, according to the decoder’s modelled action distribution.
+Each curve corresponds to a different training checkpoint of a single run.
+For both tasks, we observe that the average MSE decreases over the course of training, while the average probability of the taken actions increases.
+For later training checkpoints, we observe that the average MSE actually decreases within an episode, while the average probability increases.
+This reflects the increased confidence of the agent modelling module, as more data is observed about teammates.
+Thus, we conclude that POAM is able to cope with the challenges introduced by the sampled teammates and partial observability, to learn accurate teamate models.
+Impact of data from non-controlled agents
+Figure 5: Learning curves of POAM and IPPO, where the value network is trained with and without uncontrolled agents’ data (UCD).
+Recall that both POAM and IPPO-NAHT update the value network using data from both the controlled and uncontrolled agents.
+As Figures 5 and 9 show, this algorithmic feature results in a significant performance gain over training using on-policy data only, for both POAM and IPPO-NAHT.
+6.4 Out of Distribution Generalization
+Figure 6: Mean and standard error of test returns achieved by POAM and IPPO-NAHT, when paired with out-of-distribution teammates. POAM shows improved generalization to OOD teammates.
+The AHT literature commonly assumes that some subset of the uncontrolled teammates of interest may be interacted directly with during training [28, 6, 30].
+Experiments in the previous sections were conducted under this assumption.
+This section conducts an analysis of how POAM performs when this assumption is relaxed—that is, when the teammates have not been seen previously (Q3).
+The out-of-distribution (OOD) teammate setting is of interest because in realistic scenarios, it may be challenging to enumerate all teammates likely to be encountered in the wild.
+Here, OOD teammates are generated by using MARL algorithms with different random seeds to generate teammates that have not been interacted with during training.
+Figures 6 and 10 shows the mean and standard error of the test return achieved by POAM, compared to IPPO-NAHT, when the algorithm in question is paired with previously unseen seeds of IPPO, IQL, MAPPO, QMIX, and VDN In three out of five tasks (mpe, 5v6, 3s5z), POAM has a significantly higher return than IPPO-NAHT, while the remaining two tasks exhibit a smaller improvement.
+For all tasks, we observe that the std.
+error for POAM is lower than that of IPPO-NAHT, showing that POAM performs more consistently against a range of teammates and team configurations than IPPO-NAHT.
+7 Related Works
+This section summarizes related work in the areas most closely related to NAHT, namely, ad hoc teamwork, zero-shot coordination, agent modeling, and cooperative MARL.
+Ad Hoc Teamwork & Zero-Shot Coordination.
+Prior works in ad hoc teamwork [37] and zero-shot coordination (ZSC) [15] explored methods to design adaptive agents that can optimally collaborate with unknown teammates.
+While they both highly resemble the NAHT problem, existing methods for AHT [30, 26] and ZSC [15, 24] have been limited to single-agent control scenarios.
+We argue that direct, naive applications of AHT and ZSC techniques to our problem of interest are ineffective—see the discussion in Section 4 and results in Section 6.
+Recent research in AHT and ZSC utilizes neural networks to improve agent collaboration within various team configurations.
+These recent works mostly focus on two approaches.
+The first approach trains the agent to adapt to unknown teammates by characterizing teammates’ behavior as fixed-length vectors using neural networks and learning a policy network conditioned on these vectors [30, 28, 45].
+The second designs teammate policies that maximize the agent’s performance when collaborating with diverse teammates [24, 30].
+Our work builds on the first category, extending it to control multiple agents amid the existence of unknown teammates.
+While this also offers a potential path for robust NAHT agents, designing teammate policies for training will be kept as future work.
+Evaluating Agents’ Cooperative Capabilities.
+Beyond training agents to collaborate with teammates having unknown policies, researchers have developed environments and metrics to assess cooperative abilities.
+The Melting Pot suite [19, 1] mostly evaluates controlled agents’ ability to maximize utilitarian welfare against unknown agents.
+However, this evaluation suite focuses on mixed-motive games where agents may have conflicting goals.
+This contrasts with our work’s scope of fully cooperative settings, where all agents share the same reward function.
+MacAlpine et al.
+[25] also explored alternative metrics to measure agents’ cooperative capabilities while disentangling the effects of their overall skills in drop-in RoboSoccer.
+Agent Modeling.
+Agent modeling [2] helps us characterize other agents based on their actions.
+Such characterizations could attempt to infer important attributes of another agent’s decision making process such as their goals [12] or policies [28].
+POAM and most existing AHT techniques rely on agent modeling to provide important teammate information for decision-making when collaborating with unknown teammates.
+Cooperative MARL.
+Cooperative Multi-Agent Reinforcement Learning (MARL) explores algorithms for training agent teams on fully cooperative tasks.
+Some existing methods focus on credit assignment and decentralized control [11, 32].
+Other works in cooperative MARL also leverage parameter sharing and role assignment (e.g., [9, 42]) to decide an optimal division of labor between agents.
+However, these techniques assume control over all existing agents during training and evaluation.
+This limits these techniques’ effectiveness in settings with new teammates (NAHT problem) as shown in prior work [41, 15, 30].
+8 Conclusion
+This paper proposes and formulates the problem of $N$-agent ad hoc teamwork (NAHT), a generalization of both AHT and MARL.
+It further proposes a multi-agent reinforcement learning algorithm to train NAHT agents called POAM, and develops a procedure to train and evaluate NAHT agents.
+POAM is a policy gradient approach that uses an encoder-decoder architecture to perform teammate modeling, and leverages data from uncontrolled agents for policy optimization.
+Empirical validation on MPE and StarCraft II tasks shows that POAM consistently improves over baseline methods that naively apply existing MARL and AHT approaches, in terms of sample efficiency, asymptotic return, and generalization to out-of-distribution teammates.
+Limitations and Future Work.
+This paper addresses a special case of the NAHT problem, with homogeneous and non-communicating agents.
+POAM, which employs full parameter sharing, may not perform well in settings with heterogeneous agents, or in settings that require highly differentiated roles.
+POAM also does not leverage centralized state information or allow communication between controlled agents.
+Incorporating this information might enable learning improved NAHT policies.
+Further, POAM’s actor update is purely on-policy, and therefore cannot leverage data generated by uncontrolled agents.
+Future work may consider off-policy learning methods as a solution strategy.
+In addition to the directions suggested by POAM’s limitations, algorithmic ideas from AHT, such as diversity-based teammate generation [24] and teammate-model-based planning methods [6], also suggest rich avenues for future work.
+Having introduced the NAHT problem in this work, we hope the community explores the many potential directions to design even better NAHT algorithms by considering advances in MARL, AHT and agent modeling.
+Appendix A Appendix
+A.1 An Example of the Team Sampling Procedure
 Let $g(k,\mu,s)$ represent a function for randomly selecting $k$ elements from a generic set $\mu$, conditioned on a state $s\in\mathcal{S}$.
@@
 Returning to our robot warehouse example, a state-dependent $g$ might enable us to model that robots suitable for heavy loads are more likely to be present near the loading dock area.
+Similarly, if the time is included in the state, we would be able to model dynamic characteristics, e.g.
+that certain types of robots are only available during regular working hours, when humans are available to supervise.
+Of course, in simple cases, $g$ might be state-independent.
 For $s\in\mathcal{S}$, $X$ is a sampling procedure parameterized by the tuple $(s,M,U,C,\bm{\phi}_{M},g_{U},g_{C})$, where $\bm{\phi}_{M}\in[0,1]^{M}$ represents the logits of a categorical distribution:
@@
     Sample $M-N$ uncontrolled agents via $g_{C}$ from $C$.
-Since teams consisting entirely of unknown agents are not of interest, we restrict consideration to teams containing at least one controlled agent, i.e $N\geq 1$.
-The agent sampling functions $g_{U},g_{C}$ could employ sampling with replacement to model scenarios where two robots may use the same policy, or sampling without replacement for scenarios where that is not possible (e.g.
+Note that the agent sampling functions $g_{U},g_{C}$ could employ sampling with replacement to model scenarios where two robots may use the same policy, or sampling without replacement for scenarios where that is not possible (e.g.
 if all robots are heterogeneous).
-Without loss of generality, let $C(\theta)$ denote a set of agents parameterized by $\theta$.
-Let the notation $\bm{\pi}^{(M)}\sim X(U,C)$ indicate sampling a team of size $M$ from $U$ and $C$.
-The objective of the NAHT problem is to find parameters $\theta$, such that $C(\theta)$ maximizes the expected return in the presence of teammates from $U$:
-$\max_{\theta}\left(\mathbb{E}_{\bm{\pi}^{(M)}\sim X(U,C(\theta))}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}\right]\right).$ (1)
-The central challenges of the NAHT problem include (1) coordinating with potentially unknown teammates (generalization), and (2) coping with an unknown number of uncontrolled teammates (openness).
-4 A Motivating Example
-Having introduced the NAHT problem, a natural question to consider is whether AHT solutions may optimally address NAHT problems as well.
-If so, then there would be little need to consider the NAHT problem setting.
-For instance, as an AHT solution consists of a single teammate policy, trained in the scenario illustrated in Figure 1, a simple yet reasonable approach consists of directly copying the AHT policy as many times as required, in an NAHT problem.
-To illustrate the limitations of such an approach, we consider a concrete example in which a single ad hoc agent would follow a rational policy, but two ad hoc agents following that same policy would lead to a suboptimal outcome.
-Define the following simple game for $M$ agents: at each turn, each agent $a_{i}$ picks one bit $b_{i}\in\{0,1\}$; at the end of each turn, all the bits are summed $s=\sum_{i}b_{i}$.
-The team wins if and only if the sum of the chosen bits is exactly $1$.
-Consider a scenario where there are $M-1$ (uncontrolled) agents who follow a deterministic policy.
-For simplicity, suppose they always pick 00.
-We now introduce a new ad hoc agent $a_{h}$.
-Assume that the ad hoc agent observes the $M-1$ always-zero agents play for one step, and the ad hoc agent $a_{h}$ is introduced into the game in the second step.
-Suppose the ad hoc agent selects actions according to the following strategy:
-  1. 1.
-    Model teammate actions by observing the last step.
-  2. 2.
-    Select the best response action.
-  3. 3.
-    If no information is known about a teammate, assume they act according to the average of expected actions of all the other observed agents.
-If the ad hoc agent $a_{h}$ acts based on the teammate modeling scheme described above, it will observe that all the other agents picked 00 at the first step, and therefore pick $1$.
-Now, consider a scenario in which there are $M-2$ always-zero agents playing without the ad hoc agent for one step, and the ad hoc agents $a_{h}^{1}$ and $a_{h}^{2}$ are introduced into the game in the second step.
-Following the same modeling policy, the two ad hoc agents $a_{h}^{1}$ and $a_{h}^{2}$ will both assume that their teammates follow an always-zero policy, and therefore learn to pick the 1 action—a suboptimal decision, since $s=2$.
-Following the second step, both $a_{h}^{1}$ and $a_{h}^{2}$ will assume that their counterparts pick the $1$ action.
-As a result, both agents will switch to the action they would consider the best response, and would both pick the 00 action, meaning once again the joint action is not optimal, as $s=0$.
-We observe that extending the agent modeling window beyond one step does not resolve this inherent pattern (see Appendix A.4).
-While the policy above leads to optimal behavior for a single ad hoc agent, it results in poor joint performance in the presence of another ad hoc agent.
-A policy that explicitly accounts for the potentially unpredictable behavior of other agents and strives to cooperate with other agents would lead to better performance.
-In the example above, finding such a policy is as simple as introducing $\epsilon$-greedy exploration to the ad hoc agents’ action selection behavior.
-With that modification, both agents will continue exploring until the optimal reward of $s=1$ is achieved, in which case both agents will converge to an optimal behavior.
-This is because with high probability, the two ad hoc agents will learn to optimally coordinate by following a joint policy that selects one of the optimal joint actions of $(0,1)$ or $(1,0)$ at some time step, after which both agents will repeat that joint action due to the single-step agent modeling.
-Figure 2: The training protocol for the proposed method, POAM.
-POAM adopts an independent policy gradient architecture with parameter sharing.
-POAM trains a single policy network, which characterizes the behavior of all controlled agents.
-The yellow agents are controlled by policies drawn from $U$, the set of uncontrolled agents, while the green agents are controlled by the same POAM policy.
-Data from both controlled and uncontrolled agents is used to train the POAM critic, $V_{\pi}$, while the actor $\pi(\theta)$ is trained on data from the controlled agents only.
-The actor and critic are both conditioned on a learned team encoding vector, learned via an encoder-decoder architecture.
-5 Policy Optimization with Agent Modeling (POAM)
-This section describes the proposed Policy Optimization with Agent Modeling (POAM) method, which trains a collection of NAHT agents that can adaptively deal with different collections of unknown teammates.
-POAM relies on an agent modeling network to initially build an encoding vector of the collection of teammates encountered during an interaction.
-Adaptive agent policies that can maximize the controlled agents’ returns are then learned by POAM by training a policy network, which receives environment observation and team encoding vector as input.
-Details of the training process for both agent modeling and policy networks are described in Sections 5.1 and 5.2 respectively, while an illustration of how POAM trains NAHT agents is provided in Figure 2.
-5.1 Agent Modeling Network
-Designing adaptive policies that enable NAHT agents to achieve optimal returns against any team of uncontrolled agents drawn from some set $U$, requires information on the encountered team’s unknown behavior.
-However, in the absence of prior knowledge about uncontrolled teammates’ policies, local observations from a single timestep may not contain sufficient information regarding the encountered team.
-To circumvent this lack of information, POAM’s agent modeling network plays a crucial role in providing team encoding vectors that characterize the behavior of teammates in the encountered team.
-We identify two main criteria for desirable team encoding vectors.
-First, team encoding vectors should identify information regarding the unknown state and behavior of other agents in the environment.
-Second, team encoding vectors should ideally be computable from the sequence of local observations and actions of the team.
-Fulfilling both requirements provides an agent with useful information for decision-making in NAHT problems, even under partial observability.
-POAM produces desirable team encoding vectors by training a model with an encoder-decoder architecture, which is illustrated by red components in Figure 2.
-The encoder, $f^{\text{enc}}_{\theta^{e}}:H^{i}\mapsto\mathbb{R}^{n}$, is parameterized by $\theta^{e}$ and processes the modeling agent’s history of local observations and actions up to timestep $t$, $h^{t}=\{o_{i}^{k},a_{i}^{k-1}\}_{k=1}^{t}$, to compute a team encoding vector, $e^{t}_{i}\in\mathbb{R}^{n}$.
-This reliance on local observations helps ensure that the agent modeling network can operate without having access to the environment state (i.e.
-under partial observability).
-We then train decoder networks, $f^{\text{dec}}_{\theta^{o}}:\mathbb{R}^{n}\mapsto O^{-i}$ and $f^{\text{dec}}_{\theta^{a}}:\mathbb{R}^{n}\mapsto\Delta(A_{-i})$, that predicts the observations and actions of the modelled agents at timestep $t$, $(o_{-i}^{t},a_{-i}^{t})$, to encourage $e^{t}_{i}$ to contain relevant information for the NAHT agents’ decision-making process.
-Formally, this decoder is trained to minimize the following loss function to accurately predict the modelled agents’ observations and actions:
-$L_{\theta^{e},\theta^{o},\theta^{a}}(h^{t})=\Big{(}||f^{\text{dec}}_{\theta^{o}}(f^{\text{enc}}_{\theta^{e}}(h^{t}))-o_{-i}^{t}||^{2}-\text{log}(p(a^{t}_{-i};f^{\text{dec}}_{\theta^{a}}(f^{\text{enc}}_{\theta^{e}}(h^{t}))))\Big{)}.$ (2)
-5.2 Policy and Value Networks
-POAM relies on an actor-critic architecture to train agent policies, where the policy and networks are both conditioned on the teammate embedding produced by the encoder-decoder architecture described in Section 5.1.
-The policy network, $\pi^{i}_{\theta^{p}}:H^{i}\times\mathbb{R}^{n}\mapsto\Delta(A_{i})$, uses the NAHT agent’s local observation, $o^{t}_{i}$, and the team embedding from the encoder network, $e^{t}_{i}$, to compute a policy followed by the NAHT agents.
-Conditioning the policy network on $e^{t}_{i}$ allows an NAHT agent to change its behaviour based on the inferred characteristics of encountered agents.
-When training the policy network, we also rely on a value network $V_{\theta^{c}}:H^{i}\times\mathbb{R}^{n}\mapsto\mathbb{R}$, which measures the expected returns given $h_{t}$, and $e^{t}_{i}$.
-The value network serves as a baseline to reduce the variance of the gradient updates and also conditions on the learned teammate embeddings, for similar reasons to the policy.
-POAM then trains the policy and value networks using an approach based off the Independent PPO algorithm (Yu et al., 2022) (IPPO).
-IPPO is selected as the base MARL algorithm for two reasons.
-First, using an independent MARL method circumvents the need to deal with the changing number of agents resulting from environment openness.
-Second, it has been demonstrated to be effective on various MARL tasks.
-POAM trains the value network to produce accurate state value estimates by minimizing the following loss function:
-$L_{\theta^{c}}(h^{t})=\dfrac{1}{2}\Big{(}V_{\theta^{c}}(h^{t},f^{\text{enc}}_{\theta^{e}}(h^{t}))-\hat{V}^{t}\Big{)}^{2},$ (3)
-where $\hat{V}^{t}=\sum_{l=0}^{T}\gamma^{l}r^{t+l}$. The policy network is trained to minimize the PPO loss function, where the Generalized Advantage Estimator is employed (Schulman et al., 2015):
-$L_{\theta^{p}}(h^{t})=\text{min}\left(r^{t}(\theta^{p})\hat{A}^{t}_{GAE},\text{clip}\left(r^{t}(\theta^{p}),1-\epsilon,1+\epsilon\right)\hat{A}^{t}_{GAE}\right)-\alpha{H}(\pi^{i}_{\theta^{p}}(.|h^{t},f^{\text{enc}}_{\theta^{e}}(h^{t}))),$ (4)
-where $r^{t}(\theta^{p})=\dfrac{\pi^{i}_{\theta^{p}}(a^{t}_{i}|h^{t},f^{\text{enc}}_{\theta^{e}}(h^{t}))}{\pi^{i}_{\theta^{p_{old}}}(a^{t}_{i}|h^{t},f^{\text{enc}}_{\theta^{e}}(h^{t}))}$.
-Leveraging data from uncontrolled agents
-In the NAHT setting, we assume access to the joint observations and actions generated by the current team deployed in the environment, where the team consists of a mix of controlled and uncontrolled agents.
-POAM leverages the data from both controlled and uncontrolled agents to train the value network—in effect, treating the uncontrolled agents as exploration policies.
-This is a significant departure from PPO, which is a fully on-policy algorithm.
-However, only data from the controlled agents is used to train the policy network, as the policy update is highly sensitive to off-policy data.
-6 Experiments and Results
-This section presents an empirical evaluation of POAM and baseline approaches across different NAHT problems.
-We specifically investigate two questions.
-First, whether POAM leads to improved solutions for NAHT over baselines, in terms of sample efficiency or asymptoptic task return.
-Second, whether agent modeling enables generalization to previously unseen teammates.
-We start this section by introducing the main experimental procedures and baselines in Section 6.1.
-Section 6.2 presents our main result that compares POAM and the baselines in NAHT problems situated across different environments.
-Sections 6.36.4 then provide additional analysis on the team encoding vectors produced by POAM and its ability to generalize the learned policy to previously unseen sets of teammate policies.
-Finally, see the Appendix for further results, hyperparameters, and implementation details.
-Our code will be made available upon publication.
-6.1 Experimental Design
-This subsection describes the experimental design, including the particular NAHT instance considered in the experiments, training procedure, experimental domain, and baselines.
-6.1.1 A Practical Instantiation of NAHT
+A.2 Further Discussion of the Motivating Example
+This section provides proofs for claims in Section 4, and further discussion in support of the motivating example.
+Lemma A.1.
+For a team of $M$ agents independently and identically selecting actions with probability $p$, the $p$ that maximizes the probability of winning is $p=\frac{1}{M}$.
+Proof.
+The probability of winning, $P(s=1)$ may be computed as follows:
+$P(s=1)=M*P(\text{agent }i\text{ chooses 1 and $i^{-}$ choose 0})=Mp(1-p)^{M-1}.$
+We wish to analytically compute the global maxima of this expression with respect to $p$, by considering the boundary points $p\in{0,1}$ and the zeros of the first derivative.
+Clearly, if all agents select 1 with probability 0, the team cannot win; thus, $p=0$ is not a maxima.
+Similarly, if all agents select 1 with probability 1, the team also never wins; thus, $p=1$ is also not a maxima.
+Next, we compute the derivative of $P(s=1)$:
+$\frac{d[P(s=1)]}{dp}$ $=M[(1-p)^{M-1}-(M-1)p(1-p)^{M-2}]$
+$=M[(1-p)^{M-2}[(1-p)-(M-1)p]].$
+The zeros of the above expression occur at $p=0$ and $p=\frac{1}{M}$. For $p=\frac{1}{M}$, note that $P(s=1)=(\frac{M-1}{M})^{M-1}>0$; thus, $p=\frac{1}{M}$ must be the global maximum. ∎
+Lemma A.2.
+In a team of three agents consisting of two uncontrolled agents who select 1 with probability $p=\frac{1}{3}$, and one controlled ad hoc agent who selects 1 with probability $p_{\text{aht}}$, the probability of winning is $P(s=1)=\frac{4}{9}$.
+Proof.
+Let $s_{u}$ denote the sum of the bits chosen by the two uncontrolled agents, and $b_{aht}$ denote the bit value chosen by the controlled agent.
+The probability of winning can be computed by partitioning the winning outcomes by whether $s_{u}=1$ and the ad hoc agent selects 0, or whether the ad hoc agent selects 1 and both uncontrolled agents select 0, $s_{u}=0$.
+$P(s=1)$ $=P(s_{u}=1\land b_{aht}=0)+P(s_{u}=0\land b_{aht}=1)$
+$=2\cdot\frac{1}{3}\cdot\frac{2}{3}\cdot\bigl{(}1-p_{\text{aht}}\bigr{)}+p_{\text{aht}}\left(\frac{2}{3}\cdot\frac{2}{3}\right)$
+$=\frac{4}{9}-\frac{4}{9}\cdot p_{\text{aht}}+\frac{4}{9}\cdot p_{\text{aht}}$
+$=\frac{4}{9}.$
+Lemma A.3.
+In a team of three agents where two ad hoc agents select 1 with probability $p_{\text{aht}}$ and the remaining uncontrolled agent selects 1 with probability $\frac{1}{3}$, the maximizing $p_{\text{aht}}=\frac{1}{3}$ and the corresponding winning probability is $P(s=1)=\frac{4}{9}$.
+Proof.
+Let $s_{aht}$ denote the sum of the bits chosen by the two ad hoc agents, and $b_{u}$ denote the bit value chosen by the uncontrolled agent.
+The probability of winning may be directly computed as follows:
+$P(s=1)$ $=P(s_{aht}=1\land a_{u}=0)+P(s_{aht}=0\land a_{u}=1)$
+$=2\cdot p_{\text{aht}}\cdot(1-p_{\text{aht}})\cdot\frac{2}{3}+(1-p_{\text{aht}})^{2}\cdot\frac{1}{3}$
+$=(1-p_{\text{aht}})\left[\frac{4}{3}p_{\text{aht}}+\frac{1}{3}(1-p_{\text{aht}})\right]$
+$=(1-p_{\text{aht}})\left(\frac{1}{3}+p_{\text{aht}}\right).$
+To determine the maximizing value, we compute $P(s=1)$ at the boundary points $p_{\text{aht}}\in\{0,1\}$ and at the zeros of the derivative of $P(s=1)$.
+Note that for $p_{\text{aht}}=0$, $P(s=1)=\frac{1}{3}$, while for $p_{\text{aht}}=1$, $P(s=1)=0$.
+The derivative of the analytic expression of $P(s=1)$ with respect to $p_{\text{aht}}$ is:
+$\frac{d[P(s=1)]}{dp_{\text{aht}}}=(1-p_{\text{aht}})-\left(\frac{1}{3}+p_{\text{aht}}\right)=\frac{2}{3}-2p_{\text{aht}}.$ (4)
+The zeros of the above expression occur at $p_{\text{aht}}=\frac{1}{3}$, which corresponds to $P(s=1)=\frac{4}{9}$. ∎
+Lemma A.4.
+In a team of three agents where one agent always selects 0, one agent selects 1 with probability $\frac{1}{3}$, and the last agent selects 1 with probability $p_{\text{naht}}$, the optimal $p_{\text{naht}}=1$ and results in a winning probability of $P(s=1)=\frac{2}{3}$.
+Proof.
+Since one of the two controlled agents always plays 0, the game effectively becomes a two-player game instead.
+Let $a_{u}$ denote the action selected by the uncontrolled agent, and let $a_{naht}$ denote the action of the controlled agent. The probability of winning may be directly computed as follows:
+$P(s=1)$ $=P(a_{u}=1\land a_{naht}=0)+P(a_{u}=0\land a_{naht}=1)$
+$=\frac{1}{3}\cdot(1-p_{\text{naht}})+\frac{2}{3}\cdot p_{\text{naht}}$
+$=\frac{1}{3}+\frac{1}{3}\cdot p_{\text{naht}}.$
+It is clear that $p_{\text{naht}}=1$ maximizes $P(s=1)$, and the corresponding value of $P(s=1)=\frac{2}{3}$. ∎
+A.3 Experiment Details
+A.3.1 Team Sampling Procedure Used in Experiments
 Figure 7: A practical instantiation of the NAHT problem.
-Similar to AHT, the NAHT problem can be parameterized by the amount of knowledge that controlled agents have about uncontrolled agents, and whether uncontrolled agents can adapt to the behavior of controlled agents (Barrett & Stone, 2012).
-This section considers a practical instantiation of the NAHT problem, designed to assess whether the baseline algorithms have successfully learned to cooperate with each of the uncontrolled agents under consideration.
 Recall that $U$ and $C$ denote the sets of uncontrolled and controlled agent policies, respectively, while $M$ denotes the total team size, and $N$ the number of agents sampled from $C$.
-For simplicity, we assume that policies in $U$ do not adapt to the controlled agents’ behavior, agents are homogeneous (so that they are interchangeable within a team), and that agents cannot communicate.
-While in principle, the team sampling procedure $X$ could define any probability distribution over teams constructed from $U$ and $C$, we consider an $X_{train}$ consisting of sampling $N$ uniformly from $\{1,\cdots,M-1\}$, sampling $N$ agents from $C$ and $M-N$ agents from $U$ in a uniform fashion (Figure 3).
-The sampling procedure described here takes place at the beginning of each episode to select a team, which is then deployed in the environment.
+The experiments in the following consider an $X_{train}$ consisting of sampling $N$ uniformly from $\{1,\cdots,M-1\}$, sampling $N$ agents from $C$ and $M-N$ agents from $U$ in a uniform fashion (Figure 7).
+The sampling procedure takes place at the beginning of each episode to select a team, which is then deployed in the environment.
 Data generated by the deployed team (e.g.
@@
 See Appendix A.3.3 for further details on the evaluation procedure.
-6.1.2 Generating Uncontrolled Teammates
-To generate a set of uncontrolled teammates $U$, the following MARL algorithms are used to train agent teams: VDN (Sunehag et al., 2018), QMIX (Rashid et al., 2018), IQL (Tan, 1997), IPPO, and MAPPO (Yu et al., 2022).
-We verify that the generated team behaviors are diverse by checking that (1) teams trained by the same algorithm learn non-compatible coordination conventions, and (2) teams trained by different algorithms are not compatible (see Appendix A.2).
-Let $U_{train}$ denote the set of uncontrolled teammates used to train POAM and the baseline methods.
-$U_{train}$ consists of five teams, where each individual team is trained via VDN, QMIX, IQL, IPPO and MAPPO, respectively.
-$U_{test}$ consists of a set of holdout teams trained via the same MARL algorithms, but that have not been seen during training.
-The experimental results reported in 6.2 and 6.3 are computed with respect to $U_{train}$ only, while the experimental results in 6.4 use $U_{test}$.
-6.1.3 Experimental Domain
-Experiments are conducted on the StarCraft Multi-Agent Challenge (SMAC) benchmark (Samvelyan et al., 2019).
-SMAC features a set of cooperative tasks, where a team of allied agents must defeat a team of enemy agents controlled by the game server.
-It is a partially observable domain with a continuous state space, where each agent can observe features about itself, enemy, and allied agents within some radius.
-The action space is discrete—each agent can choose an enemy to attack, a direction to move in, or to not perform any action.
-The SMAC tasks considered here are 5v6, 8v9, 10v11, 3s4z, 3s5z.
-The first number indicates the number of allied agents, while the second indicates the number of enemy agents; full task names and descriptions are provided in App.
-A.1.
-6.1.4 Baselines
-As NAHT is a new problem proposed by this paper, there are no prior algorithms that are directly designed for the NAHT problem.
-Therefore, we construct three baselines to contextualize the performance of POAM.
-  • Naive MARL: various well-known MARL algorithms are considered, including both independent and centralized training with decentralized execution algorithms (Hernandez-Leal et al., 2019).
-The algorithms evaluated here include IQL (Claus & Boutilier, 1998), VDN (Sunehag et al., 2018), QMIX Rashid et al.
-(2018), IPPO, and MAPPO (Yu et al., 2022).
-The MARL baselines are trained in self-play and then evaluated in the NAHT setting.
-In the following, we report only the performance of the best naive MARL baseline.
-  • Independent PPO in the NAHT setting (IPPO-NAHT): IPPO is a policy gradient MARL algorithm that directly generalizes PPO (Schulman et al., 2017) to the multi-agent setting.
-It was found to be surpringly effective on a variety of MARL benchmarks by Yu et al.
-(2022).
-In contrast to the naive MARL baselines, IPPO-NAHT is trained using the proposed NAHT training scheme.
-The variant considered here employs full parameter sharing, where the actor is trained on data only from controlled agents, but the critic is trained using data from both controlled and uncontrolled agents.
-The latter detail is a key algorithmic feature which POAM also employs, but is an extension from the most naive version of PPO (see Section 6.3).
-IPPO can be considered an ablation of POAM, where the agent modeling module is removed.
-  • POAM in the AHT setting (POAM-AHT): As considered in Section 4, a natural baseline approach to the NAHT problem is to deploy AHT policies directly in the NAHT setting.
-To experimentally evaluate the intuition that AHT policies do not suffice for the NAHT problem setting, we consider an AHT version of POAM.
-POAM-AHT is trained identically to POAM, but where the number of controlled agents $N$ is always equal to 1 during training.
-All methods employ full parameter sharing (Papoudakis et al., 2021b). See Appendix A.1 for details about architecture, hyperparameter tuning and algorithm implementation.
-6.2 Main Results
-Figure 4: POAM consistently improves over the baselines of IPPO-NAHT, POAM-AHT, and the best naive MARL baseline in all tasks, in either sample efficiency or asymptotic return.
-The performance of POAM-AHT and naive MARL are inconsistent across tasks.
-Results are computed over three trials, while the shaded region represents the 95% confidence interval.
-The primary question addressed here is whether POAM improves performance on the NAHT problem setting, as compared to baselines.
-Figure 4 shows the learning curves of POAM and IPPO-NAHT on the SMAC tasks, and the test returns achieved by the best naive MARL baseline and POAM-AHT.
-See the Appendix for tables providing the performances of all naive MARL methods, across all tasks.
-All learning curves consist of the mean test returns across three trials, while the shaded regions reflect the 95% confidence intervals.
-All methods are trained for 20M time steps.
-We find that for all tasks, POAM outperforms baselines in either asymptotic return (5v6, 3s5z), or sample efficiency (8v9, 10v11).
-While POAM initially has lower sample efficiency than IPPO-NAHT, this trend is reversed after the first few million steps, after which POAM displays higher return.
-We attribute this to the initial costs incurred by learning team encoding vectors, which once learned, improves learning efficiency.
-IPPO-NAHT, which can be viewed as an ablation of POAM with no agent modeling, is the next best performing method.
-Although IPPO-NAHT has generally poorer sample efficiency than POAM, the method converges to the approximately the same returns on 8v9 and 10v11.
-While the agent modeling module of POAM provides team embedding vectors to the policy learning process, the embeddings are themselves produced from each agent’s own observation.
-Since no additional information is provided to POAM agents, it is unsurprising that IPPO-NAHT can converge to similar solutions as POAM, given enough training steps.
-Figure 5: Evolution of the within-episode mean squared error (left) and probability of actions that were actually taken by modelled teammates (right), computed by the POAM agent over the course of training.
-The agent modeling performance improves over the course of an episode, as more data about teammate behavior is observed.
-Finally, while the best naive MARL baseline and POAM-AHT learn good solutions on some tasks, neither method consistently performs well across all tasks.
-Overall, POAM discovers the most consistently performant policies compared to the other methods.
-We conclude that (1) the proposed learning scheme can learn effectively in the presence of multiple types of agents, and (2) agent modeling improves learning efficiency.
-6.3 A Closer Look at POAM
-Two key aspects of POAM are the use of a recurrent teammate modeling module, and the use of data from uncontrolled agents to train the critic.
-We perform further analysis of both of these aspects on the 5v6 task.
-Teammate modeling performance
-In the NAHT training/evaluation setting, a new set of teammates is sampled at the beginning of each episode.
-Therefore, an important subtask for a competent NAHT agent is to rapidly model the distribution and type of teammates at the beginning of the episode, to enable the policy to exploit that knowledge as the episode progresses.
-This task is especially challenging in the presence of partial observability (a property of the SMAC tasks).
-To address the above challenges, POAM employs a recurrent encoder, which encodes the POAM agent’s history of observations and actions to an embedding vector, and a (non-recurrent) decoder network, which predicts the egocentric observations and actions for all teammates.
-A natural question then, is whether the learned teammate encoding vectors actually improve over the course of an episode.
-Figure 5 depicts the within-episode mean squared error (MSE) of the observation predicted by the decoder, and the within-episode probability of the action actually taken by the modelled teammates, according to the decoder’s modelled action distribution.
-Each curve corresponds to a different training checkpoint.
-We observe the MSE rapidly converges to a low value, while the predicted probability of the taken actions has converged to 1 by time step 20.
-Note that the MSE is initially 0 at timestep 0 because the initial position is not randomized in SMAC.
-Next, we observe that the modeling performance improves over the course of training, indicated by lower prediction error at later training steps.
-Thus, we conclude that POAM learns accurate teammate models relatively early in the episode for the 5v6 task—successfully coping with the challenges introduced by the sampled teammates and partial observability.
-Impact of data from non-controlled agents
-Figure 6: Learning curves of POAM and IPPO, where the value network is trained with and without uncontrolled agents’ data (UCD).
-Recall that both POAM and IPPO-NAHT update the value network using data from both the controlled and uncontrolled agents.
-As Figure 6 shows, this algorithmic feature results in a significant performance gain over training using on-policy data only, for both POAM and IPPO-NAHT.
-6.4 Out of Distribution Generalization
-Figure 7: Mean and standard error of test returns achieved by POAM and IPPO-NAHT, when paired with out-of-distribution teammates.
-POAM shows improved generalization to OOD teammates, as compared to IPPO-NAHT, across all tasks.
-Experiments in prior sections were conducted with in-distribution teammates, i.e.
-teammates that the methods interacted directly with during training.
-This section examines whether agent modeling improves generalization to out-of-distribution (OOD) teammates, where the OOD teammates are generated by using MARL algorithms with different random seeds to generate teammates that have not been interacted with during training.
-[^4] We demonstrate in App.
-A.2 that in SMAC tasks, the same MARL algorithm with different seeds trains teams that have different coordination conventions.
-Hence, teams from unseen seeds display OOD behavior.
-Figure 7 shows the mean and standard error of the test return achieved by POAM, compared to IPPO-NAHT, when the algorithm in question is paired with previously unseen seeds of IPPO, IQL, MAPPO, QMIX, and VDN (computed across three trials).
-In two out of four tasks (5v6, 3s5z), POAM has a significantly higher return than IPPO-NAHT, while the remaining two tasks exhibit a smaller improvement.
-For all tasks, we observe that the standard error for POAM is lower than that of IPPO-NAHT, showing that POAM performs more consistently against a range of teammates and team configurations than IPPO-NAHT.
-7 Related Works
-This section summarizes related work in the areas most closely related to NAHT, namely, ad hoc teamwork, zero-shot coordination, agent modeling, and cooperative MARL.
-Ad Hoc Teamwork & Zero-Shot Coordination.
-Prior works in ad hoc teamwork (Stone et al., 2010) and zero-shot coordination (ZSC) (Hu et al., 2020) explored methods to design adaptive agents that can optimally collaborate with unknown teammates.
-While they both highly resemble the NAHT problem, existing methods for AHT (Rahman et al., 2021; Mirsky et al., 2022) and ZSC (Hu et al., 2020; Lupu et al., 2021) have been limited to single-agent control scenarios.
-We argue that direct, naive applications of AHT and ZSC techniques to our problem of interest are ineffective—see the discussion in Section 4 and results in Section 6.
-Recent research in AHT and ZSC utilizes neural networks to improve agent collaboration within various team configurations.
-These recent works mostly focus on two approaches.
-The first approach trains the agent to adapt to unknown teammates by characterizing teammates’ behavior as fixed-length vectors using neural networks and learning a policy network conditioned on these vectors (Rahman et al., 2021; Papoudakis et al., 2021a; Zintgraf et al., 2021).
-The second designs teammate policies that maximize the agent’s performance when collaborating with diverse teammates (Lupu et al., 2021; Rahman et al., 2021).
-Our work builds on the first category, extending it to control multiple agents.
-While this also offers a potential path for robust NAHT agents, designing teammate policies for training will be kept as future work.
-Evaluating Agents’ Cooperative Capabilities.
-Beyond training agents to collaborate with teammates having unknown policies, researchers have developed environments and metrics to assess cooperative abilities.
-The Melting Pot suite (Leibo et al., 2021; Agapiou et al., 2022) mostly evaluates controlled agents’ ability to maximize utilitarian welfare against unknown agents.
-However, this evaluation suite focuses on mixed-motive games where agents may have conflicting goals.
-This contrasts with our work’s scope of fully cooperative settings, where all agents share the same reward function.
-MacAlpine & Stone (2017) also explored alternative metrics to measure agents’ cooperative capabilities while disentangling the effects of their overall skills in drop-in RoboSoccer.
-Agent Modeling.
-Agent modeling (Albrecht & Stone, 2018) helps us characterize other agents based on their actions.
-Such characterizations could attempt to infer important attributes of another agent’s decision making process such as their goals (Hanna et al., 2021) or policies (Papoudakis et al., 2021a).
-POAM and most existing AHT techniques rely on agent modeling to provide important teammate information for decision-making when collaborating with unknown teammates.
-Cooperative MARL.
-Cooperative Multi-Agent Reinforcement Learning (MARL) explores algorithms for training agent teams on fully cooperative tasks.
-Some existing methods focus on credit assignment and decentralized control (Foerster et al., 2018; Rashid et al., 2018).
-Other works in cooperative MARL also leverage parameter sharing and role assignment (e.g., Christianos et al.
-(2021); Yang et al.
-(2022)) to decide an optimal division of labor between agents.
-However, these techniques assume control over all existing agents during training and evaluation.
-This limits these techniques’ effectiveness in settings with new teammates (NAHT problem) as shown in prior work (Vezhnevets et al., 2020; Hu et al., 2020; Rahman et al., 2021).
-8 Conclusion
-This paper proposes and formulates the problem of $N$-agent ad-hoc teamwork (NAHT), a generalization of both AHT and MARL.
-It further proposes a multi-agent reinforcement learning algorithm to train NAHT agents called POAM, and develops a procedure to train and evaluate NAHT agents.
-POAM is a policy gradient approach that uses an encoder-decoder architecture to perform teammate modeling, and leverages data from uncontrolled agents for policy optimization.
-Empirical validation on StarCraft II tasks shows that POAM consistently improves over baseline NAHT approaches, in terms of sample efficiency, asymptotic return, and generalization to out-of-distribution teammates.
-While this work explores teammate modeling for NAHT, other algorithmic ideas from AHT provide a rich set of future directions.
-These include generating teammates for the NAHT setting, alternative teammate modeling schemes, and considering communication protocols for NAHT agents.
-Limitations of POAM also suggest directions for future work.
-For instance, POAM is an independent MARL method (albeit with parameter sharing)—how might one leverage knowledge from the global state to generate improved NAHT policies?
-Further, POAM’s actor update is purely on-policy, and therefore cannot leverage data generated by uncontrolled agents.
-Future work may consider off-policy learning methods as a solution strategy.
-Acknowledgments
-This work has taken place in the Learning Agents Research Group (LARG) at UT Austin.
-LARG research is supported in part by NSF (FAIN-2019844, NRT-2125858), ONR (N00014-18-2243), ARO (E2061621), Bosch, Lockheed Martin, and UT Austin’s Good Systems grand challenge.
-Peter Stone serves as the Executive Director of Sony AI America and receives financial compensation for this work.
-The terms of this arrangement have been reviewed and approved by the University of Texas at Austin in accordance with its policy on objectivity in research.
-Appendix A Appendix
-A.1 Experiment Details
-A.1.1 SMAC Tasks
+A.3.2 Experimental Domains
+MPE Predator Prey
+The MPE environment [23] (released under the MIT License) is a setting where particle agents interact within a bounded 2d plane, equipped with a discrete action space and continuous observation space.
+The observation space is 16-dimensional, and contains the agents’ own position and velocity, relative positions/velocities of all other agents, landmarks, and prey.
+The action space consists of 5 discrete actions, corresponding to the four cardinal movement directions and a no-op action.
+The predator-prey task (mpe-pp) is a custom task implemented by the authors of this paper within the fork of the MPE environment released by Papoudakis et al.
+[29] (MIT License).
+In this task, three predators must cooperatively pursue a pre-trained prey agent.
+The team receives a reward of 1 if two or more predators collide with the prey at a single time step, and no reward if only a single agent collides with the prey.
+A shaping reward consisting of 0.01 times the $\ell_{2}$ distance between each agents in the team and the prey is provided as well.
+The maximum episode length is 100 time steps.
+SMAC
+SMAC [33] (released under the MIT License) features a set of cooperative tasks, where a team of allied agents must defeat a team of enemy agents controlled by the game server.
+It is a partially observable domain with a continuous state space and discrete action space.
+For each agent, the observation space is continuous, and consists of features about itself, enemy, and allied agents within some radius.
+The action space is discrete, and allows an agent to choose an enemy to attack, a direction to move in, or to not perform any action.
+As the number of allies and enemies varies between tasks, the dimensionality of the observation and action spaces is particular to each task.
+The maximum number of time steps per episode is also task specific, although early termination occurs if all enemies are defeated.
+At each time step, the team receives a shaped reward corresponding to the damage dealt, and bonuses of 10 and 200 points for killing an enemy and winning the scenario by defeating all enemies.
+The reward is scaled such that the maximal return achievable in each task is 20.
 The SMAC tasks considered in this paper are described in more detail below:
@@
   • 3s5z: stands for 3s vs 5z, three allied Stalkers versus five enemy Zealots.
-A.1.2 Evaluation Details
+Figure 8: Evolution of the within-episode mean squared error (left) and probability of actions that were actually taken by modelled teammates (right), computed by the POAM agent over the course of training, on the 5v6 task.
+The agent modeling performance improves over the course of an episode, as more data about teammate behavior is observed.
+Figure 9: Learning curves of POAM and IPPO, where the value network is trained with and without uncontrolled agents’ data (UCD) on 5v6.
+Figure 10: Mean and standard error of test returns achieved by POAM and IPPO-NAHT, when paired with out-of-distribution teammates.
+POAM shows improved generalization to OOD teammates, as compared to IPPO-NAHT, across all tasks.
+A.3.3 Evaluation Details
 The experiments in the main paper consider a particular team sampling procedure called $X_{train}$, which uniformly samples $N$ and uniformly samples agents from the sets $U$ and $C$.
@@
 The reason for this is that we are interested in the performance of agents when paired with known teammates.
-This is the model self-play score, in contrast to the algorithm self-play score (Albrecht et al., 2024).
-A.1.3 Algorithm Implementation
-The experiments in this paper use algorithm implementations from the ePyMARL codebase (Papoudakis et al., 2021b).
+This is the model self-play score, in contrast to the algorithm self-play score [3].
+A.3.4 Algorithm Implementation
+The experiments in this paper use algorithm implementations from the ePyMARL codebase [29] (released under the Apache License).
 The value-based methods are used without modification (i.e.
-IQL, VDN, QMIX), but we implement our own version of policy gradient methods (IPPO, MAPPO), based on the implementation of (Yu et al., 2022).
+IQL, VDN, QMIX), but we implement our own version of policy gradient methods (IPPO, MAPPO), based on the implementation of [43].
 All methods employ recurrent actors and critics, with full parameter sharing, i.e.
 all agents are controlled by the same policy, where the agent id is input to the policy/critic networks, to allow behavioral differentiation between agents.
+All methods employ the Adam optimizer to optimize all networks involved.
 For policy gradient methods, the policy architecture is two fully connected layers, followed by an RNN (GRU) layer, followed by an output layer.
@@
 The critic architecture is the same as the policy architecture.
-The value-based networks employ the same architecture, except that there is only a single fully connected layer before the RNN layers, and layer normalization is not used (following the ePyMARL implementation).
+The value-based networks employ the same architecture, except that there is only a single fully connected layer before the RNN layers, and layer normalization is not used, following the ePyMARL implementation.
 Please consult the code base for full implementation details.
@@
 Table 1: Hyperparameters evaluated for the policy gradient algorithms. Selected values are bolded
-Algorithm ED epochs ED Minibatches ED LR
-POAM 1, 5, 10 1, 2 0.0005, 0.005
+Task Algorithm ED epochs ED Minibatches ED LR
+5v6 POAM 1, 5, 10 1, 2 0.0005, 0.005
+mpe-pp POAM 1, 5 1, 2 0.0005, 0.001
 Table 2: Additional hyperparameters evaluated for POAM; note that ED stands for encoder-decoder. Selected values are bolded.
-A.2 Supplemental Results
-A.2.1 On coordination conventions in StarCraft II
-The experimental procedure detailed in Section 6.1 generates diverse teammates in SMAC by using MARL algorithms to train multiple teams of agents.
+A.4 Supplemental Results
+A.4.1 Validating the Existence of Coordination Conventions
+The experimental procedure detailed in Section 6.1 generates diverse teammates in SMAC and MPE by using MARL algorithms to train multiple teams of agents.
 Two underlying assumptions of the procedure are that (1) teams trained by the same algorithm learn non-compatible coordination conventions, and (2) teams trained by different algorithms are not compatible.
@@
 Self-play with MARL algorithms.
-For the SMAC tasks under consideration, teams trained using the same MARL algorithm, but different seeds, can converge to distinct coordination conventions.
+For the tasks under consideration, teams trained using the same MARL algorithm, but different seeds, can converge to distinct coordination conventions.
 Figure 11 demonstrates this by depicting the return of teams trained together (matched seeds) versus those of teams that were not trained together (mismatched seeds), across all naive MARL algorithms (IPPO, IQL, MAPPO, QMIX, VDN) and tasks considered.
 Overall, the returns of the teams that were trained together are higher than those not trained together.
-The general phenomenon has been previously observed and exploited by prior works in ad hoc teamwork (Strouse et al., 2022).
+The general phenomenon has been previously observed and exploited by prior works in ad hoc teamwork [38].
 This paper takes advantage of this to generate a set of diverse teammates for the StarCraft II experiments.
 We select tasks where the effect is significant, to ensure that there are distinct coordination behaviors for POAM to model.
-Tasks that were considered but subsequently ruled out for this reason include 3m vs 3m, 8m vs 8m, and 6h vs 8z.
+Tasks that were considered but subsequently ruled out for this reason include 3m vs 3m, 8m vs 8m, 6h vs 8z, and the MPE Spread task.
 Figure 11: Agent teams that were trained together using the same algorithm (matched seeds) have higher returns than teams that were not trained together but were trained with the same algorithm (mismatched seeds).
 Cross-play with MARL algorithms
-Tables 9, 10, 11, and 12 display the full cross-play results for all MARL algorithms on all tasks.
+Tables 12, 13, 14, 15, and 16 display the full cross-play results for all MARL algorithms on all tasks.
 Note that the tables are reflected across the diagonal axis for viewing ease.
 The values on the off-diagonal (i.e., where the two algorithms are not the same) are the cross-play score, while the values shown on the diagonal are self-play scores, computed as described in App.
-A.1.2.
+A.3.3.
 The cross-play and self-play scores displayed are means and standard errors.
@@
 We note that there are a few exceptions (e.g., IPPO vs QMIX on 8m), and that the cross-play and self-play scores are much closer on 10v11, but overall the trend is consistent across tasks.
+Figure 12: Cross-play table for the mpe-pp task.
 Figure 13: Cross-play table for the 5v6 task.
@@
 Figure 16: Cross-play table for the 10v11 task.
-A.3 Cross-play performance as a function of the number of agents.
+A.5 Cross-play performance as a function of the number of agents.
 To provide more insight on how a non-NAHT team performance changes as the number of unseen agents increases, Figure 17 plots QMIX and IPPO’s test scores on 5v6 when paired with various methods, as a function of the number of unseen agents.
@@
 The unseen agents are generated by the second method in the legend name.
-A.4 Further Discussion of the Motivating Example
-This section considers an extension of the motivating example discussed in Section 4; namely, whether considering longer agent modeling windows is sufficient to allow the ad hoc agents to solve the $s=1$ game without coordination.
-Consider a modification of the ad hoc agents’ policies such that instead of picking the best response to teammate actions in the last step, they consider the last $t$ actions of each teammate, and pick the best response with respect to the expected actions.
-In this extended scenario, assume there are $M-1$ (uncontrolled) agents which always pick 00.
-Assume that the $M-1$ always-zero agents play without the ad hoc agent for $t$ steps.
-Then, the ad hoc agent $a_{h}$ is introduced into the game in the $t+1$-th step.
-In this scenario, $a_{h}$ will learn that all the other agents always pick zero, and will pick the optimal action $1$ accordingly.
-Now consider the setting where there are $M-2$ always-zero agents playing without the ad hoc agent for $t$ steps, and two ad hoc agents $a_{h}^{1}$ and $a_{h}^{2}$ are introduced into the game at the $t+1$-th step.
-Following the same modeling policy, the two ad hoc agents $a_{h}^{1}$ and $a_{h}^{2}$ will once again both learn that their teammates follow an always-zero policy, and therefore learn to pick the $1$ action — a suboptimal decision, as $s=2$.
-In this setting, even if $a_{h}^{1}$ and $a_{h}^{2}$ modeled teammate behavior on a sliding window of length $t$ rather than every $t$ steps in batch form, they will not converge to optimal behavior.
-To see why, consider the actions of $a_{h}^{1}$ and $a_{h}^{2}$ in an arbitrary window of size $w$.
-Since they begin in unison, are deterministic, and consider the same window size, their actions will be identical for all the $t$ steps of this window.
-For the first $t$ steps, lacking history, both agents will pick action $1$, as we have shown above.
-After $k$ steps, they will both switch to 00, as we have also shown.
-Now, however, if they are both modeling agent action distributions using a rolling window of the same size $t$, after $t/2+1$ additional steps, they will reach a point in which both agents have chosen 00 more than 50% of the time in the last $t$ timesteps.
-At that point, both agents will once again simultaneously switch their actions to $1$ to maximize the likelihood of winning the game.
-As a result $S$ would once again flip from 00 to $2$.
-Using a longer window for modeling action does not mitigate the problem of inadequately modeling other ad hoc agents, which leads to an inability to cooperate.
-A.5 Computing Infrastructure
+A.6 Computing Infrastructure
 All value-based algorithms (e.g.
-QMIX, VDN, IQL) were run without parallelized training, while policy gradient algorithms were run with parallelized training.
+QMIX, VDN, IQL) were run without parallelizing environments, while policy gradient algorithms were run with parallelized environments.
+All methods were trained for 20M steps on all tasks.
+Each run took between 12-48 hours of compute, and used less than 2gB of GPU memory.
+Runs were parallelized to use computational resources most efficiently.
 The servers used in our experiments ran Ubuntu 18.04 with the following configurations:

### v3

@@
 This paper formalizes the problem, and proposes the Policy Optimization with Agent Modelling (POAM) algorithm.
-POAM is a policy gradient, multi-agent reinforcement learning approach to the NAHT problem, that enables adaptation to diverse teammate behaviors by learning representations of teammate behaviors.
+POAM is a policy gradient, multi-agent reinforcement learning approach to the NAHT problem that enables adaptation to diverse teammate behaviors by learning representations of teammate behaviors.
 Empirical evaluation on tasks from the multi-agent particle environment and StarCraft II shows that POAM improves cooperative task returns compared to baseline approaches, and enables out-of-distribution generalization to unseen teammates.
@@
 A more explicit definition of $X$ is provided in Appendix A.1.
-Without loss of generality, let $C(\theta)=\{\pi^{i}_{\theta}(.|s)\}_{i=1}^{N}$ denote a set of $N$ controlled agent policies parameterized by $\theta$, such that a learning algorithm might optimize for $\theta$.
+Without loss of generality, let $C(\theta)=\{\pi_{i}^{\theta}(.|s)\}_{i=1}^{N}$ denote a set of $M$ controlled agent policies parameterized by $\theta$, such that a learning algorithm might optimize for $\theta$.
 Let the $\bm{\pi}^{(M)}$ indicate a team of $M$ agents and $\bm{\pi}^{(M)}\sim X(U,C)$ indicate sampling such a team from $U$ and $C$ via the team sampling procedure.
@@
 Suppose the uncontrolled agents follow a policy that independently selects $1$ with probability $p=\frac{1}{M}$.
-[^3] Lemma A.1 shows that this is the optimal value of $p$ for a team composed of such agents.
+[^3] Lemma A.1 shows that $p=\frac{1}{M}$ is the optimal value of $p$ for a team composed of such agents.
 In the following, we consider the three agent case, $M=3$, for simplicity.
@@
 Let $p_{\text{aht}}$ denote the probability with which the AHT agent selects 1.
-Given the aforementioned team of uncontrolled agents, we show that any value of $p_{\text{aht}}$ results in the same $P(s=1)$, i.e.
-all possible policies for the AHT agent are optimal.
-This occurs because the probability of winning, $P(s=1)=\frac{4}{9}$, is independent of $p_{\text{aht}}$ (Lemma A.2).
+Given the aforementioned team of uncontrolled agents, we show that any value of $p_{\text{aht}}$ results in the same probability of winning, which occurs because the probability of winning, $P(s=1)=\frac{4}{9}$, is independent of $p_{\text{aht}}$ (Lemma A.2).
 Next, consider an NAHT scenario where a learning algorithm must define the actions of two out of three agents.
@@
 However, in the $N=2$ setting, we show that the optimal AHT policy $p_{\text{aht}}=\frac{1}{3}$ and the winning probability for this policy is $P(s=1)=\frac{4}{9}$ (Lemma A.3).
-Finally, we show that there exists an NAHT policy that controls both agents and obtains a higher winning probability.
+Finally, we show there exists an NAHT policy that controls both agents and obtains a higher winning probability.
 Consider the policy where one controlled agent always plays 0, while the other plays 1 with probability $p_{\text{naht}}$.
-We show that the optimal $p_{\text{naht}}=1$, and the probability of winning $P(s=1)=\frac{2}{3}>\frac{4}{9}$ (Lemma A.4).
+Lemma A.4 shows that the optimal $p_{\text{naht}}=1$, and the probability of winning $P(s=1)=\frac{2}{3}>\frac{4}{9}$.
 Thus, we have exhibited an NAHT scenario where an AHT policy that is optimal when $N=1$, performs worse than a simple NAHT joint policy in the $N=2$ setting.
+Empirical validation of the prior results are provided in Appendix A.5.1.
 5 Policy Optimization with Agent Modeling (POAM)
-Figure 2: POAM trains a single policy network, which characterizes the behavior of all controlled agents (green), while uncontrolled agents (yellow) are drawn from $U$.
-Data from both controlled and uncontrolled agents is used to train the critic, $V_{\pi}$ while the actor $\pi(\theta)$ is trained on data from the controlled agents only.
-The actor and critic are both conditioned on a learned team encoding vector.
-This section describes the proposed Policy Optimization with Agent Modeling (POAM) method to trains a collection of NAHT agents that can adaptively deal with different collections of unknown teammates.
-POAM relies on an agent modeling network to initially build an encoding vector of the collection of teammates encountered during an interaction.
-Adaptive agent policies that can maximize the controlled agents’ returns are then learned by training a policy conditioned on the environment observation and team encoding vector.
+Figure 2: POAM trains a single policy network $\pi^{\theta_{p}}$, which characterizes the behavior of all controlled agents (green), while uncontrolled agents (yellow) are drawn from $U$.
+Data from both controlled and uncontrolled agents is used to train the value network, $V^{\theta_{c}}_{i}$ while the policy is trained on data from the controlled agents only.
+The policy and value function are both conditioned on a learned team embedding vector, $e^{t}_{i}$.
+This section describes the proposed Policy Optimization with Agent Modeling (POAM) method, which trains a collection of NAHT agents that can adaptively deal with different collections of unknown teammates.
+POAM relies on an agent modeling network to initially build an embedding vector characterizing teammates encountered during an interaction.
+Adaptive agent policies that can maximize the controlled agents’ returns are then learned by training a policy conditioned on the environment observation and team embedding vector.
+To enable controlling a varying number of agents while learning in a sample-efficient manner, POAM adopts the independent learning framework with full parameter sharing.
 The training processes for agent modeling and policy networks are described in Sections 5.1 and 5.2 respectively, while an illustration of how POAM trains NAHT agents is provided in Figure 2.
@@
 However, in the absence of prior knowledge about uncontrolled teammates’ policies, local observations from a single timestep may not contain sufficient information regarding the encountered team.
-To circumvent this lack of information, POAM’s agent modeling network plays a crucial role in providing team encoding vectors that characterize the observed behavior of teammates in the encountered team.
-We identify two main criteria for desirable team encoding vectors.
-First, team encoding vectors should identify information regarding the unknown state and behavior of other agents in the environment (both controlled and uncontrolled).
-Second, team encoding vectors should ideally be computable from the sequence of local observations and actions of the team.
+To circumvent this lack of information, POAM’s agent modeling network plays a crucial role in providing team embedding vectors that characterize the observed behavior of teammates in the encountered team.
+We identify two main criteria for desirable team embedding vectors.
+First, team embedding vectors should identify information regarding the unknown state and behavior of other agents in the environment (both controlled and uncontrolled).
+Second, team embedding vectors should ideally be computable from the sequence of local observations and actions of the team.
 Fulfilling both requirements provides an agent with useful information for decision-making in NAHT problems, even under partial observability.
-POAM produces desirable team encoding vectors by training a model with an encoder-decoder architecture, which is illustrated by red components in Figure 2.
-The encoder, $f^{\text{enc}}_{\theta^{e}}:H^{i}\mapsto\mathbb{R}^{n}$, is parameterized by $\theta^{e}$ and processes the modeling agent’s history of local observations and actions up to timestep $t$, $h^{t}=\{o_{i}^{k},a_{i}^{k-1}\}_{k=1}^{t}$, to compute a team encoding vector, $e^{t}_{i}\in\mathbb{R}^{n}$.
+For each controlled agent, POAM produces informative team embedding vectors by training a model with an encoder-decoder architecture, illustrated by red components in Figure 2.
+For ease of presentation, the encoder-decoder models for controlled agent $i$ will be referred to without the index $i$.
+The encoder, $f^{\text{enc}}_{\theta^{e}}:H_{i}\mapsto\mathbb{R}^{n}$, is parameterized by $\theta^{e}$ and processes the modeling agent’s history of local observations and actions up to timestep $t$, $h_{i}^{t}=\{o_{i}^{k},a_{i}^{k-1}\}_{k=1}^{t}$, to compute a team embedding vector of dimension $n$, $e^{t}_{i}\in\mathbb{R}^{n}$ that characterizes the modeled agents.
 This reliance on local observations helps ensure that the agent modeling network can operate without having access to the environment state that is unavailable under partial observability.
-The decoder networks, $f^{\text{dec}}_{\theta^{o}}:\mathbb{R}^{n}\mapsto O^{-i}$ and $f^{\text{dec}}_{\theta^{a}}:\mathbb{R}^{n}\mapsto\Delta(A_{-i})$, are trained to predict the observations and actions of the modeled agents at timestep $t$, $(o_{-i}^{t},a_{-i}^{t})$, to encourage $e^{t}_{i}$ to contain relevant information for the NAHT agents’ decision-making process.
-Formally, this decoder is trained to minimize the following loss function to accurately predict the modeled agents’ observations and actions:
-$L_{\theta^{e},\theta^{o},\theta^{a}}(h^{t})=\Big{(}||f^{\text{dec}}_{\theta^{o}}(f^{\text{enc}}_{\theta^{e}}(h^{t}))-o_{-i}^{t}||^{2}-\text{log}(p(a^{t}_{-i};f^{\text{dec}}_{\theta^{a}}(f^{\text{enc}}_{\theta^{e}}(h^{t}))))\Big{)}.$ (2)
+The team embedding vector is decoded by two decoder networks: the observation decoder, $f^{\text{dec}}_{\theta^{o}}:\mathbb{R}^{n}\mapsto O_{-i}$ and the action decoder, $f^{\text{dec}}_{\theta^{a}}:\mathbb{R}^{n}\mapsto\Delta(A_{-i})$.
+The decoder networks are respectively trained to predict the observations and actions of all other agents on the team at timestep $t$, $(o_{-i}^{t},a_{-i}^{t})$, to encourage $e^{t}_{i}$ to contain relevant information for the current NAHT agent’s decision-making process.
+While the observation decoder directly predicts the observed $-i$ observations, the action decoder predicts the parameters of a probability distribution over the $-i$ agents’ actions, $p(a^{t}_{-i};f^{\text{dec}}_{\theta^{a}}(f^{\text{enc}}_{\theta^{e}}(h_{i}^{t})))$, where an appropriate distribution for $p$ should be selected by the system designer.
+Concretely, agent $i$’s encoder-decoder model is trained to minimize a maximum likelihood loss over all teammates’ observations and actions, given its own local observations and actions.
+As the experimental setting in this paper considers continuous observations and discrete actions, the observation loss is a mean squared error loss, while the action loss is the negative log likelihood of the $-i$ agents’ actions, under the Categorical distribution.
+$L_{\theta^{e},\theta^{o},\theta^{a}}(h_{i}^{t},o_{-i}^{t},a_{-i}^{t})=\Big{(}||f^{\text{dec}}_{\theta^{o}}(f^{\text{enc}}_{\theta^{e}}(h_{i}^{t}))-o_{-i}^{t}||^{2}-\text{log}(p(a^{t}_{-i};f^{\text{dec}}_{\theta^{a}}(f^{\text{enc}}_{\theta^{e}}(h_{i}^{t}))))\Big{)}.$ (2)
 5.2 Policy and Value Networks
 POAM relies on an actor-critic approach to train agent policies, where the policy and critic are both conditioned on the teammate embedding described in Section 5.1.
-The policy network, $\pi^{i}_{\theta^{p}}:H^{i}\times\mathbb{R}^{n}\mapsto\Delta(A_{i})$, uses the NAHT agent’s local observation, $o^{t}_{i}$, and the team embedding from the encoder network, $e^{t}_{i}$, to compute a policy followed by the NAHT agents.
+The policy network of agent $i$, $\pi_{i}^{\theta^{p}}:H_{i}\times\mathbb{R}^{n}\mapsto\Delta(A_{i})$, is parameterized by $\theta^{p}$, and uses the NAHT agent’s local observation, $o^{t}_{i}$, and the team embedding from the encoder network, $e^{t}_{i}$, to compute a policy followed by the NAHT agents.
 Conditioning the policy network on $e^{t}_{i}$ allows an NAHT agent to change its behaviour based on the inferred characteristics of encountered agents.
-When training the policy network, we also rely on a value network $V_{\theta^{c}}:H^{i}\times\mathbb{R}^{n}\mapsto\mathbb{R}$, which measures the expected returns given $h_{t}$, and $e^{t}_{i}$.
-The value network serves as a baseline to reduce the variance of the gradient updates and also conditions on the learned teammate embeddings, for similar reasons to the policy.
-POAM then trains the policy and value networks using an approach based off the Independent PPO algorithm [43] (IPPO).
+When training the policy network, we also rely on a value (or critic) network, $V_{i}^{\theta^{c}}:H_{i}\times\mathbb{R}^{n}\mapsto\mathbb{R}$, parameterized by $\theta^{c}$, which measures the expected returns given $h_{i}^{t}$, and $e^{t}_{i}$.
+The value network serves as a baseline to reduce the variance of the gradient updates, while conditioning on the learned teammate embeddings for similar reasons to the policy.
+POAM then trains the policy and value networks using an approach based on the Independent PPO algorithm [45] (IPPO).
 IPPO is selected as the base MARL algorithm for two reasons.
 First, using an independent MARL method circumvents the need to deal with the changing number of agents resulting from environment openness.
-Second, it has been demonstrated to be effective on various MARL tasks.
+Second, IPPO has been demonstrated to be effective on various MARL tasks.
+To improve learning efficiency and enable information sharing between agents, full parameter sharing is employed for all neural networks.
 POAM trains the value network to produce accurate state value estimates by minimizing the following loss function:
-$L_{\theta^{c}}(h^{t})=\dfrac{1}{2}\Big{(}V_{\theta^{c}}(h^{t},f^{\text{enc}}_{\theta^{e}}(h^{t}))-\hat{V}^{t}\Big{)}^{2},$ (3)
-where $\hat{V}^{t}=\sum_{l=0}^{T}\gamma^{l}r^{t+l}$. The policy network is trained to minimize the PPO loss function [34].
+$L_{\theta^{c}}(h_{i}^{t})=\dfrac{1}{2}\Big{(}V_{i}^{\theta^{c}}(h_{i}^{t},f^{\text{enc}}_{\theta^{e}}(h_{i}^{t}))-\hat{V_{i}}^{t}\Big{)}^{2},$ (3)
+where $\hat{V_{i}}^{t}$ is the TD($\lambda$) return.
+The policy network is analogously trained to minimize the PPO loss function [35], but where the policy additionally conditions on the team embeddings.
 Leveraging data from uncontrolled agents
-In the NAHT setting, we assume access to the joint observations and actions generated by the current team deployed in the environment, where the team consists of a mix of controlled and uncontrolled agents.
+In the NAHT setting, we assume access to the joint observations and actions generated by the current team deployed in the environment at training time only, where the team consists of a mix of controlled and uncontrolled agents.
+This assumption provides an opportunity to learn useful cooperative behaviors more quickly, by bootstrapping based on transitions from the initially more competent, uncontrolled teammate policies, as also observed by Rahman et al.
+[31].
 POAM leverages data from both controlled and uncontrolled agents to train the value network—in effect, treating the uncontrolled agents as exploration policies.
-This aspect of POAM is a significant departure from PPO, which is a fully on-policy algorithm.
-However, only data from the controlled agents is used to train the policy network, as the policy update is highly sensitive to off-policy data.
+Note that this aspect of POAM is a significant departure from PPO, which is a fully on-policy algorithm.
+Since the policy update is highly sensitive to off-policy data, only data from the controlled agents is used to train the policy network.
 6 Experiments and Results
 This section presents an empirical evaluation of POAM and baseline approaches across different NAHT problems. We investigate three questions and foreshadow the conclusions:
-  • Q1: Does POAM lead to better sample efficiency and/or asymptotic return than baselines? (Usually)
-  • Q2: Can we verify that the two key ideas of POAM—agent modelling and use of data from uncontrolled agents—work as desired and contribute positively towards POAM’s performance? (Yes)
-  • Q3: Does agent modelling improve generalization to previously unseen teammates? (Yes)
-Full implementation details, including all hyperparameter values, along with additional empirical results, appear in Appendix.
-Our code is included with the submission and will be open-sourced upon publication.
+  • Q1: Does POAM learn to cope with uncontrolled teammates with higher sample efficiency and asymptotic return than baselines? (Usually)
+  • Q2: Does POAM improve generalization to previously unseen and out-of-distribution teammates, compared to baselines? (Yes)
+  • Q3: Can we verify that the two key ideas of POAM—agent modelling and use of data from uncontrolled agents—work as desired and contribute positively towards POAM’s performance? (Yes)
+Full implementation details, including hyperparameter values and additional empirical results, appear in the Appendix. Our code is available at https://github.com/carolinewang01/naht.
 6.1 Experimental Design
-This subsection describes the experimental design, including the particular NAHT problem setting, training procedure, experimental domain, and baselines.
+In the following, we summarize the experimental design, including the particular NAHT problem instance, training procedure, experimental domain, and baselines.
+Details on evaluation metrics are provided in Appendix A.3.3.
 A Practical Instantiation of NAHT
@@
 To generate a set of uncontrolled teammates $U$, the following MARL algorithms are used to train agent teams: VDN [41], QMIX [33], IQL [42], IPPO, and MAPPO [45].
-We verify that the generated team behaviors are diverse by checking that (1) teams trained by the same algorithm learn non-compatible coordination conventions, and (2) teams trained by different algorithms are not compatible (Appendix A.4).
-This emergence of diverse teammate behaviors from training agents using different MARL algorithms under different seeds aligns with the findings from Strouse et al.
+We verify that the generated team behaviors are diverse by checking that (1) teams trained by the same algorithm learn non-compatible coordination conventions, and (2) teams trained by different algorithms also learn non-compatible coordination conventions (Appendix A.5.2).
+The emergence of diverse teammate behaviors from training agents using different MARL algorithms under different seeds aligns with the findings from Strouse et al.
 [40].
-Let $U_{train}$ denote the set of uncontrolled teammates used to train POAM and the baseline methods.
+Let $U_{train}$ denote the set of uncontrolled teammates used to train all (N)AHT methods.
 $U_{train}$ consists of five teams, where each individual team is trained via VDN, QMIX, IQL, IPPO and MAPPO, respectively.
 $U_{test}$ consists of a set of holdout teams trained via the same MARL algorithms, but that have not been seen during training.
-The experimental results reported in 6.2 and 6.3 are computed with respect to $U_{train}$ only, while the experimental results in 6.4 use $U_{test}$.
+The experimental results reported in Sections 6.2 and 6.4 are computed with respect to $U_{train}$ only, while the experimental results in Section 6.3 use $U_{test}$.
 Experimental Domains
@@
 IPPO can be considered an ablation of POAM, where the agent modeling module is removed.
-  • POAM in the AHT setting (POAM-AHT): As considered in Section 4, a natural baseline approach to the NAHT problem is use AHT algorithms that train only a single controlled agent, and copy these policies as many times as needed in the NAHT setting.
-To evaluate the intuition that AHT policies do not suffice for the NAHT problem setting, we consider an AHT version of POAM that is trained identically to POAM, but where the number of controlled agents $N=1$ during training.
+  • POAM in the AHT setting (POAM-AHT): As considered in Section 4, a natural baseline approach to the NAHT problem is to use AHT algorithms that train only a single controlled agent, and copy these policies as many times as needed in the NAHT setting.
+To evaluate the intuition that AHT policies do not suffice for the NAHT problem setting, we consider an AHT version of POAM that is trained identically to POAM, but where the number of controlled agents is always one ($N=1$) during training.
+Note that POAM-AHT is equivalent to the AHT algorithm introduced by Papoudakis et al.
+[29], LIAM.
 6.2 Main Results
 Figure 3: POAM consistently improves over the baselines of IPPO-NAHT, POAM-AHT, and the best naive MARL baseline in all tasks, in either sample efficiency or asymptotic return.
-Results are computed over three trials, while the shaded region represents the 95% confidence interval.
-This section addresses Q1—that is, whether POAM improves asymptotic return and/or sample efficiency, compared to baselines.
+This section addresses Q1—that is, whether POAM learns to cope with uncontrolled teammates with greater sample efficiency or asymptotic returns, compared to baselines.
 Figure 3 shows the learning curves of POAM and IPPO-NAHT, and the test returns achieved by the best naive MARL baseline and POAM-AHT, on all tasks.
 [^4] See the Appendix for tables providing the performances of all naive MARL methods, across all tasks.
-All learning curves consist of the mean test returns across three trials, while the shaded regions reflect the 95% confidence intervals.
-We find that for all tasks, POAM outperforms baselines in terms of asymptotic return for three out of five tasks (pp, 5v6, 3s5z).
+All learning curves consist of the mean test returns across five trials, while the shaded regions reflect the 95% confidence intervals.
+We find that for all tasks, POAM outperforms baselines in terms of asymptotic return for three out of five tasks (mpe-pp, 5v6, 3s5z).
 For all tasks, POAM’s initial sample efficiency is similar to that of IPPO-NAHT for the first few million steps of training, after which POAM displays higher return.
-We attribute this to the initial costs incurred by learning team encoding vectors, which once learned, improves learning efficiency.
+We attribute the initial similarity in sample efficiency to the initial cost incurred by learning team embedding vectors, which once learned, improves learning efficiency.
 IPPO-NAHT, which can be viewed as an ablation of POAM with no agent modeling, is the next best performing method.
@@
 Since no additional information is provided to POAM agents, it is unsurprising that IPPO-NAHT can converge to similar solutions as POAM, given enough training steps.
-Figure 4: Evolution of the within-episode mean squared error (left) and the within-episode probability of actions of modelled teammates (right), computed by the POAM agent over the course of training on the mpe-pp task.
+Figure 4: Test returns achieved by POAM and IPPO-NAHT, when paired with out-of-distribution teammates. POAM has improved generalization to OOD teammates, compared with IPPO-NAHT.
 Finally, while the best naive MARL baseline and POAM-AHT learn good solutions on some tasks, neither method consistently performs well across all tasks.
-Overall, POAM discovers the most consistently performant policies compared to baseline methods, in a sample-efficient manner.
+Overall, POAM discovers the most consistently performant policies compared to baseline methods, in a relatively sample-efficient manner.
 We conclude that (1) agent modeling improves learning efficiency, and (2) direct duplication of AHT-trained policies is less effective than methods that co-train agents for NAHT.
+6.3 Out of Distribution Generalization
+The AHT literature commonly assumes that uncontrolled teammates of interest may be interacted directly with during training [29, 6, 31]; experiments in the prior section were conducted under this assumption.
+However, in realistic scenarios, it may be challenging to enumerate all teammates likely to be encountered in the wild.
+This section examines the effectiveness of POAM under a true NAHT evaluation scenario, where POAM agents must coordinate with teammates that were not available at training time and are out-of-distribution (OOD) (Q2).
+Here, OOD teammates are created by running MARL algorithms with different random seeds than those used to generate train-time teammates.
+Figures 4 and 10 show the mean and 95% confidence intervals of the test return achieved by POAM, compared to IPPO-NAHT, when the algorithm in question is paired with previously unseen seeds of IPPO, IQL, MAPPO, QMIX, and VDN.
+For each type of teammate, the performance of IPPO-NAHT/POAM against the exact teammates seen during training is shown as the in-distribution baseline.
+Both POAM and IPPO-NAHT consistently exhibit reduced performance against the OOD teammates, compared to their respective in-distribution performances.
+In three out of five tasks (mpe, 5v6, 3s5z), POAM has a significantly higher return than IPPO-NAHT, while the remaining two tasks exhibit a smaller improvement.
+In Appendix A.5.3, similar findings are presented with an alternative OOD teammate generation strategy, where the set of five MARL algorithms used to generate uncontrolled teammates (IPPO, IQL, MAPPO, QMIX, VDN) are divided into train/test sets.
 6.4 A Closer Look at POAM
-Two key aspects of POAM are the use of a recurrent teammate modeling module, and the use of data from uncontrolled agents to train the critic (Q2).
-We perform further analysis of both aspects on the mpe-pp and 5v6 tasks; results on 5v6 may be found in Appendix A.4.
+Two key aspects of POAM are the teammate modeling module, and the use of data from uncontrolled agents to train the critic (Q3).
+We study the impact of both aspects on POAM’s sample efficiency, focusing on the mpe-pp and 5v6 tasks.
+Results on 5v6 may be found in Appendix A.4.
 Teammate modeling performance
@@
 This task is especially challenging in the presence of partial observability (a property of the SMAC tasks).
-To address the above challenges, POAM employs a recurrent encoder, which encodes the POAM agent’s history of observations and actions to an embedding vector, and a (non-recurrent) decoder network, which predicts the egocentric observations and actions for all teammates.
-A natural question then, is whether the learned teammate encoding vectors actually improve over the course of an episode.
-Figures 4 and 9 depict the within-episode mean squared error (MSE) of the observation predicted by the decoder, and the within-episode probability of the action actually taken by the modelled teammates, according to the decoder’s modelled action distribution.
+To address the above challenges, POAM employs a recurrent encoder, which encodes the POAM agent’s history of observations and actions to an embedding vector, and a (non-recurrent) decoder network, which predicts the egocentric observations and action distribution for all teammates.
+Figure 5: Evolution of a POAM agent’s within-episode mean squared error (left) and within-episode probability of actions of modeled teammates (right), over the course of training on mpe-pp.
+A natural question then, is whether the learned teammate embedding vectors actually improve over the course of an episode.
+Figures 5 and 9 depict the within-episode mean squared error (MSE) of the observation predicted by the decoder, and the within-episode probability of the action actually taken by the modeled teammates, according to the decoder’s modeled action distribution.
 Each curve corresponds to a different training checkpoint of a single run.
 For both tasks, we observe that the average MSE decreases over the course of training, while the average probability of the taken actions increases.
-For later training checkpoints, we observe that the average MSE actually decreases within an episode, while the average probability increases.
-This reflects the increased confidence of the agent modelling module, as more data is observed about teammates.
-Thus, we conclude that POAM is able to cope with the challenges introduced by the sampled teammates and partial observability, to learn accurate teamate models.
+For later training checkpoints, we observe that the average MSE actually decreases within an episode, while the average probability increases, reflecting the increased confidence of the agent modelling module as more data is observed about teammates.
+Thus, we conclude that POAM is able to cope with the challenges introduced by the sampled teammates and partial observability, to learn accurate teammate models.
 Impact of data from non-controlled agents
-Figure 5: Learning curves of POAM and IPPO, where the value network is trained with and without uncontrolled agents’ data (UCD).
 Recall that both POAM and IPPO-NAHT update the value network using data from both the controlled and uncontrolled agents.
 As Figures 6 and 9 show, this algorithmic feature results in a significant performance gain over training using on-policy data only, for both POAM and IPPO-NAHT.
-6.4 Out of Distribution Generalization
-Figure 6: Mean and standard error of test returns achieved by POAM and IPPO-NAHT, when paired with out-of-distribution teammates. POAM shows improved generalization to OOD teammates.
-The AHT literature commonly assumes that some subset of the uncontrolled teammates of interest may be interacted directly with during training [28, 6, 30].
-Experiments in the previous sections were conducted under this assumption.
-This section conducts an analysis of how POAM performs when this assumption is relaxed—that is, when the teammates have not been seen previously (Q3).
-The out-of-distribution (OOD) teammate setting is of interest because in realistic scenarios, it may be challenging to enumerate all teammates likely to be encountered in the wild.
-Here, OOD teammates are generated by using MARL algorithms with different random seeds to generate teammates that have not been interacted with during training.
-Figures 6 and 10 shows the mean and standard error of the test return achieved by POAM, compared to IPPO-NAHT, when the algorithm in question is paired with previously unseen seeds of IPPO, IQL, MAPPO, QMIX, and VDN In three out of five tasks (mpe, 5v6, 3s5z), POAM has a significantly higher return than IPPO-NAHT, while the remaining two tasks exhibit a smaller improvement.
-For all tasks, we observe that the std.
-error for POAM is lower than that of IPPO-NAHT, showing that POAM performs more consistently against a range of teammates and team configurations than IPPO-NAHT.
+Figure 6: Learning curves of POAM and IPPO-NAHT, where the value network is trained w/w.o. uncontrolled agents’ data (UCD).
 7 Related Works
-This section summarizes related work in the areas most closely related to NAHT, namely, ad hoc teamwork, zero-shot coordination, agent modeling, and cooperative MARL.
+This section summarizes literature in areas most closely related to NAHT, namely, ad hoc teamwork, zero-shot coordination, evaluation of cooperative capabilities, agent modeling, and CMARL.
 Ad Hoc Teamwork & Zero-Shot Coordination.
@@
 Agent Modeling.
-Agent modeling [2] helps us characterize other agents based on their actions.
-Such characterizations could attempt to infer important attributes of another agent’s decision making process such as their goals [12] or policies [28].
-POAM and most existing AHT techniques rely on agent modeling to provide important teammate information for decision-making when collaborating with unknown teammates.
-Cooperative MARL.
-Cooperative Multi-Agent Reinforcement Learning (MARL) explores algorithms for training agent teams on fully cooperative tasks.
+Agent modeling enables agents to characterize other agents based on their actions [2].
+Such characterizations could attempt to directly infer modeled agents’ actions, goals [13], or policies [29].
+The modeled attributes have been used in cooperative, competitive, and general sum settings to inform update rules [11, 36] or to directly inform decision making [29].
+POAM relies on agent modeling to provide important teammate information for decision-making when collaborating with unknown teammates.
+Cooperative MARL (CMARL).
+CMARL explores algorithms for training agent teams on fully cooperative tasks.
 Some existing methods focus on credit assignment and decentralized control [12, 33].
-Other works in cooperative MARL also leverage parameter sharing and role assignment (e.g., [9, 42]) to decide an optimal division of labor between agents.
-However, these techniques assume control over all existing agents during training and evaluation.
-This limits these techniques’ effectiveness in settings with new teammates (NAHT problem) as shown in prior work [41, 15, 30].
+Other works in CMARL also leverage parameter sharing and role assignment (e.g., [9, 44]) to decide an optimal division of labor between agents.
+However, these techniques assume control over all existing agents during training and evaluation, which limits their effectiveness in settings with unseen or uncontrolled teammates, as shown in prior work [43, 16, 31].
 8 Conclusion
@@
 Further, POAM’s actor update is purely on-policy, and therefore cannot leverage data generated by uncontrolled agents.
-Future work may consider off-policy learning methods as a solution strategy.
+Future work might consider employing off-policy methods to exploit the uncontrolled agent data.
 In addition to the directions suggested by POAM’s limitations, algorithmic ideas from AHT, such as diversity-based teammate generation [25] and teammate-model-based planning methods [6], also suggest rich avenues for future work.
-Having introduced the NAHT problem in this work, we hope the community explores the many potential directions to design even better NAHT algorithms by considering advances in MARL, AHT and agent modeling.
+Having introduced the NAHT problem in this work, we hope the community explores the many potential directions to design even better NAHT algorithms by considering advances in MARL, AHT, and agent modeling.
+Acknowledgments and Disclosure of Funding
+This work has taken place in the Learning Agents Research Group (LARG) at UT Austin.
+LARG research is supported in part by NSF (FAIN-2019844, NRT-2125858), ONR (N00014-18-2243), ARO (W911NF-23-2-0004, W911NF-17-2-0181), Lockheed Martin, and UT Austin’s Good Systems grand challenge.
+Peter Stone serves as the Executive Director of Sony AI America and receives financial compensation for this work.
+The terms of this arrangement have been reviewed and approved by the University of Texas at Austin in accordance with its policy on objectivity in research.
 Appendix A Appendix
@@
 $P(s=1)=M*P(\text{agent }i\text{ chooses 1 and $i^{-}$ choose 0})=Mp(1-p)^{M-1}.$
-We wish to analytically compute the global maxima of this expression with respect to $p$, by considering the boundary points $p\in{0,1}$ and the zeros of the first derivative.
+We wish to analytically compute the global maxima of this expression with respect to $p$, by considering the boundary points $p\in\{0,1\}$ and the zeros of the first derivative.
 Clearly, if all agents select 1 with probability 0, the team cannot win; thus, $p=0$ is not a maxima.
@@
 Figure 7: A practical instantiation of the NAHT problem.
-Recall that $U$ and $C$ denote the sets of uncontrolled and controlled agent policies, respectively, while $M$ denotes the total team size, and $N$ the number of agents sampled from $C$.
+Recall that $U$ and $C$ denote the sets of uncontrolled and controlled agent policies, respectively, while $M$ denotes the team size for the task, and $N$ the number of agents sampled from $C$.
 The experiments in the following consider an $X_{train}$ consisting of sampling $N$ uniformly from $\{1,\cdots,M-1\}$, sampling $N$ agents from $C$ and $M-N$ agents from $U$ in a uniform fashion (Figure 7).
@@
 The observation space is 16-dimensional, and contains the agents’ own position and velocity, relative positions/velocities of all other agents, landmarks, and prey.
-The action space consists of 5 discrete actions, corresponding to the four cardinal movement directions and a no-op action.
+The action space consists of five discrete actions, corresponding to the four cardinal movement directions and a no-op action.
+The observation range is normalized to $[-1,1]$, while the discrete action space is one-hot encoded.
 The predator-prey task (mpe-pp) is a custom task implemented by the authors of this paper within the fork of the MPE environment released by Papoudakis et al.
@@
 In this task, three predators must cooperatively pursue a pre-trained prey agent.
+We use the pre-trained prey policy provided by the ePymarl MARL framework.
+The prey policy was originally trained by Papoudakis et al.
+[30], by using the MADDPG MARL algorithm to train both predator and prey agents for 25M steps, and generally attempts to escape approaching predators.
 The team receives a reward of 1 if two or more predators collide with the prey at a single time step, and no reward if only a single agent collides with the prey.
 A shaping reward consisting of 0.01 times the $\ell_{2}$ distance between each agents in the team and the prey is provided as well.
+Since the prey policy is pre-trained and fixed, note that our predator-prey task is a fully cooperative task from the perspective of the predator (learning) agents.
 The maximum episode length is 100 time steps.
@@
 As the number of allies and enemies varies between tasks, the dimensionality of the observation and action spaces is particular to each task.
+The observation space is normalized to be between $[-1,1]$, while the action space is one-hot encoded.
 The maximum number of time steps per episode is also task specific, although early termination occurs if all enemies are defeated.
@@
   • 3s5z: stands for 3s vs 5z, three allied Stalkers versus five enemy Zealots.
-Figure 8: Evolution of the within-episode mean squared error (left) and probability of actions that were actually taken by modelled teammates (right), computed by the POAM agent over the course of training, on the 5v6 task.
-The agent modeling performance improves over the course of an episode, as more data about teammate behavior is observed.
-Figure 9: Learning curves of POAM and IPPO, where the value network is trained with and without uncontrolled agents’ data (UCD) on 5v6.
-Figure 10: Mean and standard error of test returns achieved by POAM and IPPO-NAHT, when paired with out-of-distribution teammates.
-POAM shows improved generalization to OOD teammates, as compared to IPPO-NAHT, across all tasks.
 A.3.3 Evaluation Details
-The experiments in the main paper consider a particular team sampling procedure called $X_{train}$, which uniformly samples $N$ and uniformly samples agents from the sets $U$ and $C$.
-While the learning curves reported in the main paper are computed directly with $X_{train}$, we use an iterative and deterministic evaluation procedure for the results presented in Section 6.4 and in the Appendix.
-The purpose of the additional evaluation procedure presented here is to ensure that we thoroughly evaluate the trained NAHT agents for all possible values of $N$ and all possible teammates.
-Given a set of uncontrolled agents $U$ and a set of agents $C$ generated by some NAHT method, we compute the following $M-N$ cross-play scores.
-Let $N$ be the number of agents sampled from $C$, such that $N<M$.
-For $N\in\{1,\cdots,M-1\}$, construct the joint policy $\bm{\pi}^{(N)}$ by selecting $N$ agents uniformly from $C$ and $M-N$ agents from $U$.
+This section provides details on how mixed teams are evaluated in the NAHT setting, and how cross-play scores, and self-play scores, and uncertainty measures are computed.
+$M-N$ score.
+Given a set of controlled agents $C$, and a set of uncontrolled agents $U$, the goal of the $M-N$ score is to quantify the performance when these two teams must cooperate within the NAHT scenario.
+The $M-N$ score is computed in a deterministic and exhaustive fashion, by iterating over all possible values of $N$.
+Let $N$ be the number of agents sampled from set $C$, such that $N<M$.
+For $N\in\{1,\cdots,M-1\}$, construct the joint policy $\bm{\pi}^{(M)}$ by selecting $N$ agents uniformly from $C$ and $M-N$ agents from $U$.
 Evaluate the resultant team on the task for $E$ episodes.
-This results in $M-1$ episode returns, on which we may compute summary statistics.
-If agents in $U$ and $C$ are generated by algorithm $A$ and $B$, respectively, then we call the thus computed average return, the $M-N$ cross-play score.
-More concretely: each algorithm (VDN, QMIX, QMIX-NS, IQL, IQL-NS) is run $k$ times with different seeds, to generate $k$ teams of teammates to act as uncontrolled agents.
+This results in $(M-1)*E$ episode returns, which is averaged to form the $M-N$ score.
+Cross-play scores.
+The cross-play scores reported in this paper are the average returns of teams generated by algorithm $A$, when coordinating with those generated by $B$.
+To compute the cross-play score, we first train multiple teams (by varying the seed) via both algorithms $A$ and $B$.
+Next, the $M-N$ score is computed for random pairings of teams (seeds) from $A$ and $B$, following the procedure specified above.
+Summary statistics can be computed over the set of all such NAHT returns.
+For example: each algorithm (VDN, QMIX, IQL, MAPPO, IPPO) is run $k$ times with different seeds, to generate $k$ teams of agents that may act as uncontrolled teammates.
 For each pair of algorithms, we sample a subset of the possible seed pairs, and evaluate the teams that result from merging said seed pairs.
-Note that each seed actually represents a team of agents.
-For example, if VDN and QMIX have seeds 1 2, and 3, the cross-play evaluation might consider the seed pairings (VDN 1, QMIX 2), (VDN 2, QMIX 3), (VDN 3, QMIX 1).
+If VDN and QMIX have seeds 1 2, and 3, the cross-play evaluation might consider the seed pairings (VDN 1, QMIX 2), (VDN 2, QMIX 3), (VDN 3, QMIX 1).
 Given a pair of seeds (e.g.
 (VDN 1), (QMIX 2)), the $M-N$ cross-play score is computed as the average return generated by sweeping $N\in\{1,\cdots,M-1\}$ and evaluating the merged team that consists of selecting $N$ agents from the first team, and $M-N$ agents for $E$ episodes.
-In our experiments, $E=128$, and $k=3$.
-Means and standard errors are computed over all seed pairings considered, as we treat each $M-N$ evaluation as an independent trial.
-For most experiments, three seed pairs are considered, where seeds are paired to ensure that all seeds participate in at least one evaluation.
-This ensures that the computational cost of the evaluation remains linear in $k$, rather than quadratic.
+In our experiments, $E=128$, and $k=5$.
+Self-play scores.
+The self-play scores reported in this paper are model self-play score, rather than algorithm self-play score [3].
+The reason for this is that we are interested in the performance of agents when paired with known teammates.
+Measuring uncertainty
+Means and 95% confidence intervals are computed over all team/seed pairings considered for any given scores, as we treat each $M-N$ evaluation as an independent trial.
+For most experiments, five seed pairs are considered, where seeds are paired to ensure that all seeds participate in at least one evaluation.
+This ensures that the computational cost of the evaluation remains linear in $N$, rather than quadratic.
 However, note that for the OOD experiments presented in Section 6.3, we consider all possible seed pairings for the most comprehensive evaluation.
-Self-play scores.
-For algorithm pairs that are the same, seeds are paired with themselves.
-For example, the seed pairs for VDN versus VDN would be $(1,1),(2,2),(3,3)$.
-The reason for this is that we are interested in the performance of agents when paired with known teammates.
-This is the model self-play score, in contrast to the algorithm self-play score [3].
 A.3.4 Algorithm Implementation
@@
 The value-based methods are used without modification (i.e.
-IQL, VDN, QMIX), but we implement our own version of policy gradient methods (IPPO, MAPPO), based on the implementation of [43].
+IQL, VDN, QMIX), but we implement our version of policy gradient methods (IPPO, MAPPO), based on the implementation of Yu et al.
+[45].
+Parameter Sharing.
 All methods employ recurrent actors and critics, with full parameter sharing, i.e.
-all agents are controlled by the same policy, where the agent id is input to the policy/critic networks, to allow behavioral differentiation between agents.
-All methods employ the Adam optimizer to optimize all networks involved.
+all agents are controlled by the same policy, where the agent id is input to the actor and critic networks, to allow behavioral differentiation between agents.
+For POAM, which also maintains encoder-decoder networks for agent modeling, parameter sharing is used for the encoder and decoder networks, to improve training sample efficiency.
+Further, POAM’s decoder, which predicts observations and actions for all $-i$ agents, employs parameter sharing across agent predictions, to prevent the target dimensionality from scaling with the number of teammates.
+Optimizer and Neural Architecture.
+The Adam optimizer is applied for all networks involved.
 For policy gradient methods, the policy architecture is two fully connected layers, followed by an RNN (GRU) layer, followed by an output layer.
@@
 The critic architecture is the same as the policy architecture.
-The value-based networks employ the same architecture, except that there is only a single fully connected layer before the RNN layers, and layer normalization is not used, following the ePyMARL implementation.
+The value-based methods employ the same architecture, except that there is only a single fully connected layer before the RNN layers, and layer normalization is not used, following the ePyMARL implementation.
 Please consult the codebase for full implementation details.
+Algorithm Buffer size Epochs Minibatches Entropy Clip Clip value loss
+IPPO 128, 256, 512 1, 4, 10 1, 3 0.01, 0.03, 0.05 0.01, 0.05, 0.1 no, yes
+MAPPO 64, 128, 256 4, 10 1, 3 0.01, 0.03, 0.05, 0.07 0.05, 0.1, 0.2 no, yes
+Table 1: Hyperparameters evaluated for the policy gradient algorithms. Selected values are bolded
 Hyperparameters.
@@
 POAM adopts the same hyperparameters as IPPO where applicable.
-We also tuned additional hyperparameters specific to POAM (see Table 2).
-Algorithm Buffer size Epochs Minibatches Entropy Clip Clip value loss
-IPPO 128, 256, 512 1, 4, 10 1, 3 0.01, 0.03, 0.05 0.01, 0.05, 0.1 no, yes
-MAPPO 64, 128, 256 4, 10 1, 3 0.01, 0.03, 0.05, 0.07 0.05, 0.1, 0.2 no, yes
-Table 1: Hyperparameters evaluated for the policy gradient algorithms. Selected values are bolded
+We also tuned additional hyperparameters specific to POAM (Table 2).
 Task Algorithm ED epochs ED Minibatches ED LR
@@
 Table 2: Additional hyperparameters evaluated for POAM; note that ED stands for encoder-decoder. Selected values are bolded.
-A.4 Supplemental Results
-A.4.1 Validating the Existence of Coordination Conventions
+A.4 Supplemental Figures
+This section contains additional figures referenced by the main paper. Please see Section 6 for the corresponding analysis and discussion.
+Figure 8: Evolution of the within-episode mean squared error (left) and probability of actions that were actually taken by modeled teammates (right), computed by the POAM agent over the course of training, on the 5v6 task.
+The agent modeling performance improves over the course of an episode, as more data about teammate behavior is observed.
+Figure 9: Learning curves of POAM and IPPO, where the value network is trained with and without uncontrolled agents’ data (UCD) on 5v6.
+Figure 10: Returns achieved by POAM and IPPO-NAHT, when paired with out-of-distribution teammates.
+POAM shows improved generalization to OOD teammates, as compared to IPPO-NAHT, across all StarCraft tasks.
+For each type of teammate, the performance of IPPO-NAHT/POAM against the exact teammate seen during training is shown as the in-distribution baseline.
+A.5 Supplemental Analysis
+The following subsections contains secondary analysis and results, intended to support the primary analysis of the main paper.
+For all results, the mean and 95% confidence interval over five trials is reported.
+All reported returns are test returns.
+A.5.1 The Need for Dedicated NAHT Algorithms - Empirical Evidence
+Section 4 argues theoretically that with a population of uncontrolled teammates that select one with probability 1/3 and zero otherwise, an optimal policy learned in the AHT setting would not be optimal in the NAHT setting.
+Here, we present experiments on the three agent bit matrix game that empirically verify the theory.
+Let $N$ denote the number of controlled agents.
+The episode length is 25, and the observation for each agent consists of the agent index and the joint action at the previous time step.
+The team reward at each time step is $3*\mathbbm{1}_{\sum_{i}b_{i}=1}$.
+The optimal expected return in the $N=1$ (AHT) setting is 33.333 (derived from Lemma A.3), and in the $N=2$ (NAHT) setting is 50.0 (derived from Lemma A.4).
+The methods compared are POAM-AHT and POAM.
+Table 3 shows the expected returns achieved by POAM and POAM-AHT on the $N=1$ and $N=2$ scenarios.
+Since any policy is optimal in the $N=1$ case, as expected, both methods achieve near-optimal returns.
+In the $N=2$ case, as expected, POAM outperforms POAM-AHT by a large margin, achieving a near-optimal return of $48.858\pm 1.092$.
+N=1 (AHT) N=2 (NAHT)
+POAM-AHT 33.441 $\pm$ 0.511 22.5 $\pm$ 8.250
+POAM 33.483 $\pm$ 0.485 48.858 $\pm$ 1.092
+Table 3: Returns of POAM-AHT versus POAM on the three agent bit matrix game, in the N=1 (AHT) and N=2 (NAHT) setting.
+POAM and POAM-AHT both achieve the optimal return for the $N=1$ case, but POAM has a much higher return on the $N=2$ case.
+A.5.2 Validating the Existence of Coordination Conventions
 The experimental procedure detailed in Section 6.1 generates diverse teammates in SMAC and MPE by using MARL algorithms to train multiple teams of agents.
@@
 Cross-play with MARL algorithms
-Tables 12, 13, 14, 15, and 16 display the full cross-play results for all MARL algorithms on all tasks.
+Table 4: Cross-play results for the mpe-pp task.
+Table 5: Cross-play results for the 5v6 task.
+Table 6: Cross-play results for the 8m task.
+Table 7: Cross-play results for the 3s5z task.
+Table 8: Cross-play results for the 10v11 task.
+Tables 4, 5, 6, 7, and 8 display the full cross-play results for all MARL algorithms and POAM-AHT, on all tasks.
 Note that the tables are reflected across the diagonal axis for viewing ease.
@@
 A.3.3.
-The cross-play and self-play scores displayed are means and standard errors.
-The rightmost column reflects the row average and standard deviation of the cross-play (xp) scores corresponding to the test set of VDN, QMIX, IQL, IPPO, and MAPPO, excluding the self-play score.
-Thus, the rightmost column reflects how well on average the row algorithm can generalize to unfamiliar teams.
+The cross-play and self-play scores displayed are means and 95% confidence intervals.
+The rightmost column reflects the row average and 95% confidence interval of the cross-play (XP) scores corresponding to the test set of VDN, QMIX, IQL, IPPO, and MAPPO, excluding the self-play score.
+Thus, the rightmost column reflects how well on average the row MARL/AHT algorithm can generalize to the test set used throughout this paper.
 Overall, we find that MARL algorithms perform significantly better in self-play than cross-play.
 We note that there are a few exceptions (e.g., IPPO vs QMIX on 8m), and that the cross-play and self-play scores are much closer on 10v11, but overall the trend is consistent across tasks.
-Figure 12: Cross-play table for the mpe-pp task.
-Figure 13: Cross-play table for the 5v6 task.
-Figure 14: Cross-play table for the 8m task.
-Figure 15: Cross-play table for the 3s5z task.
-Figure 16: Cross-play table for the 10v11 task.
-A.5 Cross-play performance as a function of the number of agents.
-To provide more insight on how a non-NAHT team performance changes as the number of unseen agents increases, Figure 17 plots QMIX and IPPO’s test scores on 5v6 when paired with various methods, as a function of the number of unseen agents.
-Note that $k=0$ corresponds to teams entirely composed of the evaluated agents, while $k=5$ corresponds to teams entirely composed of the unseen teammates, and are actually self-play scores, whereas the other points are cross-play scores.
-These points appear on the plot to contextualize the performance of the mixed teams.
-QMIX and IPPO have similar returns in self-play, but comparing the two figures, we observe a difference in the rate at which the cross-play returns decline.
-In particular, QMIX’s performance declines most around $k=3$ across a variety of methods, while IPPO’s performance is minimized at $k=2$.
-Figure 17: Mean and std.
-error of QMIX (left) and IPPO (right) test returns on 5v6, in the presence of uncontrolled agents generated by various MARL learning algorithms, as the number of unseen agents increases from $k=0$ to a maximum of $k=5$.
-The unseen agents are generated by the second method in the legend name.
+A.5.3 Generalization to Unseen Teammate Types
+As discussed and experimentally verified in Section A.5.2, diverse coordination conventions in StarCraft and MPE Predator Prey can be generated by (1) running the same MARL algorithm with various random seeds, or (2) running various MARL algorithms.
+In Section 6.3, the out-of-distribution (OOD) teammates considered were generated by running MARL algorithms with different random seeds than those used to generate train-time teammates—in other words, generating diverse and unseen teammates using the first procedure specified above.
+Figure 12: Returns achieved by POAM and IPPO-NAHT, when trained on MAPPO, QMIX, and IQL, and tested on IPPO, VDN.
+Here, we generate OOD teammates using the second procedure.
+More precisely, the five MARL algorithms used to generate teammates may be divided into train/test sets.
+Figure 12 shows the results of such an experiment, where POAM and IPPO-NAHT are trained with MAPPO, QMIX, and IQL teammates, and tested with agents from IPPO and VDN.
+Similar to the results presented in Section 6.3, we find that POAM generally outperforms IPPO-NAHT for the unseen teammate types for all tasks.
+A.5.4 Modeling Controlled and Uncontrolled Agents
+Figure 13: Loss of POAM’s encoder-decoder on the mpe-pp task, separated by uncontrolled (left) and controlled agents (right).
+POAM’s encoder-decoder (ED) models both controlled and uncontrolled agents.
+Since the controlled agents are updated during the training process, the encoder-decoder must deal with a “moving target".
+Thus, in principle, the problem of modeling the controlled agents is more challenging than modeling uncontrolled agents.
+Fig.
+13 shows the probability of predicting the correct action for the uncontrolled agents and controlled agents separately, on the mpe-pp task.
+Note that the action probabilities shown in Fig.
+5 (right) of the main paper would be the average of the two plots shown in Fig.
+13.
+For uncontrolled agents (Fig.
+13, left), we observe that the accuracy of action predictions for the uncontrolled agents increases much more consistently as training goes on, and is higher than that for the controlled agents.
+As expected, the ED is able to model the uncontrolled agents more easily than the nonstationary controlled agents.
+A.5.5 Performance as a Function of the Number of Controlled Agents
+To provide more insight on how an NAHT team’s performance changes as the number of controlled agents increases, Figure 14 displays the mean test returns of POAM and POAM-AHT as a function of the number of controlled agents, where the evaluation returns are averaged across five types of uncontrolled teammates: QMIX, VDN, IQL, MAPPO, and IPPO.
+Note that the self-play returns of POAM and POAM-AHT (i.e.
+where the number of controlled agents is maximal) are shown as horizontal lines, while the performances at $N=0$ correspond to the averaged self-play returns of the five uncontrolled team types.
+Recall that POAM-AHT was trained on the $N=1$ scenario only, while POAM is trained on the full NAHT setting.
+As expected, POAM outperforms POAM-AHT for all values of $N>1$, while for $N=1$, the methods perform similarly.
+POAM-AHT’s performance declines as the number of controlled agents increases, which likely occurs because the evaluation setting becomes further from the training setting as $N$ increases.
+Figure 14: Comparing the performance of POAM versus POAM-AHT, as the number of controlled agents varies.
+POAM and POAM-AHT agents are evaluated with the following set of uncontrolled agents: QMIX, VDN, IQL, MAPPO, IPPO.
 A.6 Computing Infrastructure
@@
 Each run took between 12-48 hours of compute, and used less than 2gB of GPU memory.
-Runs were parallelized to use computational resources most efficiently.
-The servers used in our experiments ran Ubuntu 18.04 with the following configurations:
+Runs were parallelized to use computational resources efficiently.
+The servers used for our experiments ran Ubuntu 20.04 with the following configurations:
   • Intel Xeon CPU E5-2630 v4; Nvidia Titan V GPU.